### v1

# Convergence analysis of Levenberg-Marquardt method with Singular Scaling for non-zero residue nonlinear least-squares problems [^1] This work is supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]

Abstract
This paper addresses the convergence analysis of a variant of the Levenberg-Marquardt method (LMM) designed for nonlinear least-squares problems with non-zero residue. This variant, called LMM with Singular Scaling (LMMSS), allows the LMM scaling matrix to be singular, encompassing a broader class of regularizers, which has proven useful in certain applications. In order to establish local convergence results, a careful choice of the LMM parameter is made based on the gradient linearization error, dictated by the nonlinearity and size of the residual. Under completeness and local error bound assumptions we prove that the distance from an iterate to the set of stationary points goes to zero superlinearly and that the iterative sequence converges. Furthermore, we also study a globalized version of the method obtained using line-search and prove that any limit point of the generated sequence is stationary. Some examples are provided to illustrate our theoretical results.
Keywords. Levenberg-Marquardt, Singular Scaling Matrix, Convergence analysis, Non-zero residue
1 Introduction
In this paper, we investigate global and local convergence properties of the Levenberg–Marquardt method applied to a nonlinear least-squares (NLS) problem:
$\min_{x\in\mathbb{R}^{n}}\quad\frac{1}{2}\|F(x)\|^{2}:=\phi(x),$ (1)
where $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is twice continuously differentiable. In particular, we focus on the overdetermined problem, where $m\geq n$.
Unlike previous research, which predominantly focused on the zero residue case or in cases where the rank of the Jacobian is complete [7, 6, 10, 11, 13], here we consider NLS problems where the set of zeros of $F$ is empty and its Jacobian is allowed to be rank-deficient.
We say that problem (1) is a non-zero-residue NLS problem when the set of zeros of $F$ is empty. Our interest is to identify the stationary points of $\phi$, i.e., the set $X^{*}=\{x\in\mathbb{R}^{n}\mid J(x)^{T}F(x)=0\}$, where $J(x)\in\mathbb{R}^{m\times n}$ is the Jacobian of $F$ at $x$, under the assumption that $X^{*}\neq\emptyset$. Throughout the text we shall use $x^{*}$ to denote an arbitrary element of $X^{*}$.
Although works such as [2, 3] explore problems with non-zero residuals and present results of local convergence of the LM method, it should be noted that the scale matrix used in these works is the identity matrix. In this work, in addition to assuming non-zero residue, we consider the following iteration:
$(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}$ $=-J_{k}^{T}F_{k}$ (2)
$x_{k+1}$ $=x_{k}+\alpha_{k}d_{k},\quad\forall k\geq 0,$ (3)
where $F_{k}:=F(x_{k})$, $J_{k}:=J(x_{k})$ is the Jacobian of $F$ at $x_{k}$, $\lambda_{k}>0$, $\alpha_{k}$ is the step size, and $L^{T}L$ referred to as scaling matrix, is allowed to be singular. We refer to the iteration (2)–(3) as the Levenberg-Marquardt method with Singular Scaling (LMMSS).
This type of iteration is motivated by applications in inverse problems, as recently seen in [5], where a parameter recovery problem using LMMSS is addressed, including a local convergence analysis for the zero residue case. On the other hand, convergence properties of the classic LMM in the non-zero residue scenario have received recent attention in the literature [2, 3]. These are two aspects in which our contribution differs from the existing literature.
Our aim is to analyze the local convergence of LMM with singular scaling matrices in the non-zero residue scenario. Furthermore, we shall provide a globalization for the LMMSS iteration where the step-size $\alpha_{k}$ is determined by line-search and prove that any limit point of the generated sequence is stationary for the least-squares function.
This paper is organized as follows. In Section 2, we start with some assumptions that will be considered throughout the manuscript. In Section 3, we present technical results necessary for the upcoming sections. We show that LMMSS iteration is well-defined and provide the local convergence results in Section 4, which are categorized into two cases based on the behavior of the rank of the Jacobian around a stationary point (constant and diminishing rank). Finally, Section 5 proves that, for LMMSS with an Armijo line-search scheme, every limit point of the generated sequence is a stationary point of the sum-of-squares function, thereby establishing the global convergence of this method. A few examples are given in Section 6 to illustrate our theoretical results. Final considerations are given in Section 7.
2 Assumptions
Recall that throughout this manuscript, we assume that the set of stationary points of $\phi$ is non-empty: $X^{*}\neq\emptyset$.
Assumption 1 below guarantees that $d_{k}$ solution of (2) is unique.
Assumption 1.
The matrix $L\in\mathbb{R}^{p\times n}$ has $\text{rank}(L)=p\leq n\leq m$ and there exists $\Omega\subset\mathbb{R}^{n}$ and $\gamma>0$ such that, for every $x\in\Omega$
$\left\lVert J(x)v\right\rVert^{2}+\left\lVert Lv\right\rVert^{2}\geq\gamma\left\lVert v\right\rVert^{2},\quad\forall v\in\mathbb{R}^{n}.$ (4)
Such assumption comes from the inverse problems literature [12, 8] where it is sometimes called completeness condition. It was a key assumption in [5] to study the local convergence of LMMSS in the zero residue setting. We remark that such assumption is equivalent to $\mathcal{N}(J(x))\cap\mathcal{N}(L)=\{{\bf 0}\},\forall x\in\Omega.$
Next assumption is quite standard in the Gauss–Newton and Levenberg–Marquardt literature and asks the Jacobian to be locally Lipschitz.
Assumption 2.
For any $x^{*}\in X^{*}$, there exists a constant $\delta\in(0,1)$ and $L_{0}>0$ such that
$\|J(x)-J(y)\|\leq L_{0}\|x-y\|,$ (5)
for all $x,y\in B(x^{*},\delta)$.
Assumption 2 implies that:
$\|J(y)(x-y)-(F(x)-F(y))\|\leq L_{1}\|x-y\|^{2},\quad\forall x,y\in B(x^{*},\delta)$ (6)
where $L_{1}=L_{0}/2$, that is, the error in the linear approximation of $F(x)$ around $y$ is $O(\|x-y\|^{2})$, for $x$ and $y$ in a neighborhood of $x^{*}$.
Due to the compactness of the ball $B(x^{*},\delta)$, there exist positive constants $L_{2}$ and $\beta$ such that $\|J(x)\|\leq L_{2}$ and $\|F(x)\|\leq\beta$ for all $x\in B(x^{*},\delta)$. Therefore, since $\|J(x)\|$ is bounded in $B(x^{*},\delta)$, by applying the mean value inequality, we can infer that $\|F(x)-F(y)\|\leq L_{2}\|x-y\|$, for all $x,y\in B(x^{*},\delta)$. Additionally, the gradient $\nabla\phi(x)=J(x)^{T}F(x)$ is Lipschitz in $B(x^{*},\delta)$:
$\left\lVert J(x)^{T}F(x)-J(y)^{T}F(y)\right\rVert\leq L_{3}\|x-y\|,\quad\forall x,y\in B(x^{*},\delta),$ (7)
where $L_{3}=L_{2}^{2}+\beta L_{0}.$
Moreover, notice that for $z\in X^{*}\cap B(x^{*},r)$ and $x,y\in B(x^{*},r)$, we have
$\|(J(x)-J(y))^{T}F(y)\|$ $=\|(J(x)-J(z)+J(z)-J(y))^{T}F(y)\|$ (8)
$\leq\|(J(x)-J(z))^{T}F(y)\|+\|(J(z)-J(y))^{T}F(y)\|$
$\leq L_{0}L_{2}\|x-z\|\|y-z\|+\|J(x)^{T}F(z)\|$
$+L_{0}L_{2}\|y-z\|^{2}+\|J(y)^{T}F(z)\|.$
Lemma 1.
[2, Lemma 2.1] If Assumption 2 is satisfied, then there exists a constant $L_{4}>0$ such that
$\left\lVert\nabla\phi(y)-\nabla\phi(x)-J(x)^{T}J(x)(y-x)\right\rVert\leq L_{4}\left\lVert x-y\right\rVert^{2}+\left\lVert(J(x)-J(y))^{T}F(y)\right\rVert,$ (9)
for all $x,y\in B(x^{*},\delta)$.
Next, we present the local error bound condition assumed in this work. We remark that it is an error bound condition with respect to the set of stationary points of $\phi$ rather than the zeroes of $F$ (which is empty in the case of non-zero residue).
Assumption 3 (Error bound).
For any $x^{*}\in X^{*}$, $\left\lVert J(x)^{T}F(x)\right\rVert$ provides an error bound in $B(x^{*},\delta)$, i.e., there exists $\omega\in(0,\infty)$ such that
$\omega\textup{dist}(x,X^{*})\leq\left\lVert J(x)^{T}F(x)\right\rVert,\quad\forall x\in B(x^{*},\delta),$ (10)
where $\textup{dist}(x,X^{*})=\inf_{z\in X^{*}}{\left\lVert x-z\right\rVert}$.
Throughout the text, given $x$, we shall denote by $\bar{x}$ an element of $X^{*}$ such that $\|x-\bar{x}\|=\text{dist}(x,X^{*})$. Also, let $x^{*}\in X^{*}$ and $J_{*}=J(x^{*})$.
From Assumptions 3 and (7), we obtain
$\omega\text{dist}(x_{k},X^{*})\leq\left\lVert J_{k}^{T}F_{k}\right\rVert\leq L_{3}\text{dist}(x_{k},X^{*}).$ (11)
The remaining assumption focuses, as outlined in (8), on the terms
$\left\lVert J(x)^{T}F(z)\right\rVert\text{ and }\left\lVert J(y)^{T}F(z)\right\rVert.$
These terms play a crucial role in controlling the error of the “incomplete linearization” of the gradient, as expressed in (9[2].
Assumption 4.
For every $x\in B(x^{*},\delta)$ and every $z\in X^{*}\cap B(x^{*},\delta)$, the following inequality holds:
$\|(J(x)-J(z))^{T}F(z)\|\leq C\|x-z\|^{1+r},$ (12)
with $r\in(0,1]$ and $C\geq 0$.
3 Auxiliary Results
In this section, we organize other fundamental results essential for the upcoming analysis.
We start by mentioning that (2) always has a solution. In fact, (2) corresponds to the normal equation $B^{T}Bd=B^{T}c$ of the augmented system
$Bd:=\left[\begin{array}[]{c}J_{k}\\ \sqrt{\lambda_{k}}L\end{array}\right]d=-\left[\begin{array}[]{c}F_{k}\\ 0\end{array}\right]=:c,$
and the normal equation always has a solution because $B^{T}c$ lies in the range of $B^{T}B$. But it is Assumption 1 that ensures the solution of (2) is unique. Indeed, it is easy to see that condition (4) at $x_{k}$ implies that $J_{k}^{T}J_{k}+\lambda_{k}L^{T}L$ is positive definite (and thus nonsingular).
Next result will be important when the rank of $J(x)^{T}J(x)$ is constant in a neighborhood of $x^{*}$.
Lemma 2.
[2, Corollary 3.1] Given $\kappa>1$, if
$\text{rank}(J(x)^{T}J(x))=\text{rank}(J_{*}^{T}J_{*})=q\geq 1,$
and
$\left\lVert J(x)^{T}J(x)-J_{*}^{T}J_{*}\right\rVert\leq\Big{(}1-\frac{1}{\kappa}\Big{)}\frac{1}{\left\lVert(J_{*}^{T}J_{*})^{+}\right\rVert},$ (13)
where $(J_{*}^{T}J_{*})^{+}$ denote the Moore-Penrose pseudo-inverse, then
$\left\lVert(J(x)^{T}J(x))^{+}\right\rVert\leq\kappa\left\lVert(J_{*}^{T}J_{*})^{+}\right\rVert.$
Now, we will show some results on the Generalized Singular Value Decomposition (GSVD), an important tool in theoretical analysis of the LMMSS direction.
Theorem 1.
(GSVD) [9, p. 22] Consider the pair $(A,L)$, where $A\in\mathbb{R}^{m\times n}$, $L\in\mathbb{R}^{p\times n}$, $m\geq n\geq p$, $rank(L)=p$, and $\mathcal{N}(A)\cap\mathcal{N}(L)=\{0\}$. Then, there exist matrices $U\in\mathbb{R}^{m\times n}$ and $V\in\mathbb{R}^{p\times p}$ with orthonormal columns and a nonsingular matrix $X\in\mathbb{R}^{n\times n}$ such that
$A=U\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right]X^{-1}\quad\text{and}\quad L=V\left[\begin{array}[]{cc}M&0\end{array}\right]X^{-1},$ (14)
with $\Sigma$ and $M$ being the following diagonal matrices:
$\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{p})\in\mathbb{R}^{p\times p}\quad\text{and}\quad M=\text{diag}(\mu_{1},\dots,\mu_{p})\in\mathbb{R}^{p\times p}.$ (15)
Moreover, the elements of $\Sigma$ and $M$ are non-negative, ordered as follows:
$0\leq\sigma_{1}\leq\dots\leq\sigma_{p}\leq 1\quad\text{and}\quad 1\geq\mu_{1}\geq\dots\geq\mu_{p}>0,$ (16)
and normalized by the relation $\sigma_{i}^{2}+\mu_{i}^{2}=1$, for $i=1,\dots,p$. We call the generalized singular value of the pair $(A,L)$ the ratio
$\gamma_{i}=\frac{\sigma_{i}}{\mu_{i}},\quad i=1,\dots,p.$
Remark 1.
Since $X^{-1}$ is nonsingular and $U$ has orthonormal columns we have
$\text{rank}(A)=\text{rank}(U\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right]X^{-1})=\text{rank}(\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right])\geq n-p.$
By considering the GSVD of the pair $(J_{k},L)$ we can provide a useful characterization of the direction $d_{k}$ in LMMSS. In fact, given the GSVD
$J_{k}=U_{k}\left[\begin{array}[]{cc}\Sigma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}\quad\text{and}\quad L=V_{k}\left[\begin{array}[]{cc}M_{k}&0\end{array}\right]X_{k}^{-1},$
where
$(\Sigma_{k})_{ii}:=\sigma_{i,k}\quad\text{and}\quad(M_{k})_{ii}:=\mu_{i,k},\quad i=1,\dots,p;$ (17)
it follows that
$J_{k}^{T}J_{k}+\lambda_{k}L^{T}L=X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}.$ (18)
Then, $d_{k}$ from (2) can be expressed as:
$d_{k}=-X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}{J_{k}}^{T}F_{k},$ (19)
with $\Gamma_{k}:=(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}.$
Next result from [5] gives an upper bound on $\|X_{k}\|$ depending on $\gamma$ from Assumption 1.
Lemma 3.
[5, Lemma 2.2] Consider $\{x_{k}\}$ the sequence generated by LMMSS along with the GSVD of the pair $(J_{k},L)$. Then, $\forall k\in\mathbb{N}$, $\|X_{k}\|\leq 1/\sqrt{\gamma}$.
Remark 2.
By the GSVD for the pair $(J_{k},L)$, it follows that
$J_{k}^{T}J_{k}+L^{T}L=X_{k}^{-T}X_{k}^{-1}.$
Therefore,
$\left\lVert X_{k}^{-1}\right\rVert^{2}\leq\left\lVert J_{k}^{T}J_{k}\right\rVert+\left\lVert L\right\rVert^{2}\leq L_{2}^{2}+\left\lVert L\right\rVert^{2}.$ (20)
Now, let us analyze an upper bound for $\left\lVert\Gamma_{k}\right\rVert$. From the GSVD for the pair $(J_{k},L)$, using the generalized singular values $\gamma_{i,k}=\sigma_{i,k}/\mu_{i,k}$, we have:
$\sigma_{i,k}^{2}=\frac{\gamma_{i,k}^{2}}{\gamma_{i,k}^{2}+1}\quad\text{and}\quad\mu_{i,k}^{2}=\frac{1}{\gamma_{i,k}^{2}+1}.$
Thus, we can express $\Gamma_{k}$ from (19) as:
$(\Gamma_{k})_{ii}=\frac{1}{\sigma_{i,k}^{2}+\lambda_{k}\mu_{i,k}^{2}}=\frac{\gamma_{i,k}^{2}+1}{\gamma_{i,k}^{2}+\lambda_{k}},\quad i=1,\dots,p.$
Therefore, we can bound $\|\Gamma_{k}\|$ in terms of the generalized singular values $\gamma_{i,k}$ and the parameter $\lambda_{k}$. To this end, we can use the following lemma.
Lemma 4.
For the function
$\psi(\gamma,\lambda)=\dfrac{\gamma^{2}+1}{\gamma^{2}+\lambda},\quad\gamma\geq 0,\ \lambda>0,$
the following properties hold
  • (a)
    For $\lambda\in(0,1)$, the function $\psi(\gamma,\lambda)$ has a unique maximum value attained at
    $\gamma_{\max}=0\quad\text{ where }\quad\psi(\gamma_{\max},\lambda)=\max_{\gamma\geq 0}\psi(\gamma,\lambda)=\frac{1}{\lambda}.$
  • (b)
    For a fixed $\lambda\in[1,+\infty)$, the function $\psi(\gamma,\lambda)$ is non-decreasing and upper-bounded. More precisely, $\psi(\gamma,\lambda)\leq 1,\quad\forall\gamma\geq 0.$
Proof.
First, note that for any fixed $\lambda$, we have
$\lim_{\gamma\rightarrow\infty}\psi(\gamma,\lambda)=1.$ (21)
Now, given $\lambda>0$, observe that
$\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)=\frac{2(\lambda-1)\gamma}{(\gamma^{2}+\lambda)^{2}}.$ (22)
  • (a)
    Consider $\lambda\in(0,1)$ fixed. As the maximizers of $\psi(\gamma,\lambda)$, for $\gamma>0$, must satisfy $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)=0,$ from (22), we have $02(\lambda-1)\gamma=$ 0, thus $\gamma_{\max}=0$. Substituting $\gamma_{\max}$ into $\psi(\gamma,\lambda)$, we obtain
    $\psi(\gamma_{\max},\lambda)=1/2.$
  • (b)
    Now, for fixed $\lambda\in[1,+\infty)$, from (22), we have $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)>0$ for $\gamma\geq 0$. Therefore, $\psi$ is increasing for $\gamma\in[0,+\infty)$, with $\psi(0,\lambda)=0$. Then, from (21), we conclude that $\psi(\gamma,\lambda)\leq 1$, for all $\gamma\geq 0$.
We are now ready to prove two key lemmas for the local convergence analysis. They show that the norm of LMMSS direction $d_{k}$ is ${\cal O}(\textup{dist}(x_{k},X^{*}))$, considering two distinct cases. Lemma 5 addresses the scenario where the rank of the Jacobian around the solution (stationary-point) is constant, while Lemma 6 deals with the case of diminishing rank.
Lemma 5.
Suppose that Assumptions 1-3 are valid in $B(x^{*},\delta)$, and
$rank(J(x)^{T}J(x))=rank(J(x^{*})^{T}J(x^{*}))=q\geq 1$
for all $x\in B(x^{*},\delta)$. If $x_{k}\in B(x^{*},\delta)$, $\lambda_{k}>0$ and $J_{k}^{T}F_{k}\neq 0$, then there exists $c_{1}>0$ such that
$\|d_{k}\|\leq c_{1}\textup{dist}(x_{k},X^{*}).$ (23)
Proof.
Recall that $J_{k}^{T}F_{k}\in\mathcal{R}(J_{k}^{T})=\mathcal{N}(J_{k})^{\perp}=\mathcal{N}(J_{k}^{T}J_{k})^{\perp}$, and
$(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}=-J_{k}^{T}F_{k}.$ (24)
First, notice that $d_{k}\notin\mathcal{N}(J_{k})$. In fact, from (24), we have
$d_{k}^{T}(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}=-(J_{k}d_{k})^{T}F_{k},$
which implies that $\left\lVert J_{k}d_{k}\right\rVert^{2}+\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=-(J_{k}d_{k})^{T}F_{k}$, thus, if we assume that $d_{k}\in\mathcal{N}(J_{k})$, we conclude that $\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=0$. But $\left\lVert Ld_{k}\right\rVert>0$, given that, by Assumption 1, $\mathcal{N}(J_{k})\cap\mathcal{N}(L)=0$, and also $\lambda_{k}\neq 0$, and we have a contradiction.
We can write $d_{k}=d_{N}+d_{R},$ with $d_{N}\in\mathcal{N}(J_{k})$ and $d_{R}\in\mathcal{N}(J_{k})^{\perp}=\mathcal{R}(J_{k}^{T})$. Since $d_{k}\notin\mathcal{N}(J_{k})$, we have $d_{R}\neq 0$.
Thus, from (24)
$-J_{k}^{T}F_{k}$ $=(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}=(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)(d_{N}+d_{R})$ (25)
$=J_{k}^{T}J_{k}d_{N}+J_{k}^{T}J_{k}d_{R}+\lambda_{k}L^{T}Ld_{N}+\lambda_{k}L^{T}Ld_{R}$
$=J_{k}^{T}J_{k}d_{R}+\lambda_{k}L^{T}Ld_{N}+\lambda_{k}L^{T}Ld_{R}.$
When we multiply (25) by $d_{N}^{T}$, we get
$(J_{k}d_{N})^{T}J_{k}d_{R}+\lambda_{k}\left\lVert Ld_{N}\right\rVert^{2}+\lambda_{k}d_{N}^{T}L^{T}Ld_{R}=-(J_{k}d_{N})^{T}F_{k},$
and as $d_{N}\in\mathcal{N}(J_{k})$ and $\lambda_{k}>0$, we have:
$\left\lVert Ld_{N}\right\rVert^{2}=-(Ld_{N})^{T}Ld_{R}.$ (26)
From (26), we also conclude that $\left\lVert Ld_{N}\right\rVert^{2}\leq\left\lVert Ld_{N}\right\rVert\left\lVert Ld_{R}\right\rVert$, and as $d_{N}\notin{\cal N}(L)$ we obtain
$\left\lVert Ld_{N}\right\rVert\leq\left\lVert Ld_{R}\right\rVert.$ (27)
On the other hand, when we multiply (25) by $d_{R}^{T}$, we get
$\left\lVert J_{k}d_{R}\right\rVert^{2}+\lambda_{k}(Ld_{R})^{T}Ld_{N}+\lambda_{k}\left\lVert Ld_{R}\right\rVert^{2}=-(J_{k}d_{R})^{T}F_{k}.$ (28)
Using (26) in (28), we are left with
$\left\lVert J_{k}d_{R}\right\rVert^{2}+\lambda_{k}\left\lVert Ld_{R}\right\rVert^{2}=-(J_{k}d_{R})^{T}F_{k}+\lambda_{k}\left\lVert Ld_{N}\right\rVert^{2}.$
From this and (27), we have
$\left\lVert J_{k}d_{R}\right\rVert^{2}\leq-(J_{k}d_{R})^{T}F_{k}.$ (29)
Let $J_{k}=U\Sigma V^{T}$ be the SVD for $J_{k}$. As $d_{R}\in\mathcal{N}(J_{k})^{\perp}$, $d_{R}$ can be written as a linear combination of the columns of $V$ corresponding to the non-zero singular values of $J_{k}$. Therefore,
$\left\lVert J_{k}d_{R}\right\rVert=\left\lVert U\Sigma V^{T}d_{R}\right\rVert\geq s_{q,k}\left\lVert V^{T}d_{R}\right\rVert=s_{q,k}\left\lVert d_{R}\right\rVert,$
where $s_{q,k}$ is the smallest positive singular value of $J_{k}$. By using $s_{q,k}\left\lVert d_{R}\right\rVert\leq\left\lVert J_{k}d_{R}\right\rVert$ in (29):
$s_{q,k}^{2}\left\lVert d_{R}\right\rVert^{2}\leq-(J_{k}d_{R})^{T}F_{k}\leq\left\lVert J_{k}^{T}F_{k}\right\rVert\left\lVert d_{R}\right\rVert,$
which implies
$\left\lVert d_{R}\right\rVert\leq\frac{1}{s_{q,k}^{2}}\left\lVert J_{k}^{T}F_{k}\right\rVert.$ (30)
Now let us show that $d_{R}\notin\mathcal{N}(L)$. Suppose $d_{R}\in\mathcal{N}(L)$, from (26), we get $\left\lVert Ld_{N}\right\rVert^{2}=-d_{N}^{T}L^{T}Ld_{R}=0,$ implying that $d_{N}\in\mathcal{N}(L)$ and $d_{N}\in\mathcal{N}(J_{k})$, which is a contradiction with Assumption 1.
By Assumption 1, there exists $\gamma$ such that $\gamma\left\lVert d_{N}\right\rVert^{2}\leq\left\lVert J_{k}d_{N}\right\rVert^{2}+\left\lVert Ld_{N}\right\rVert^{2}$, which implies
$\left\lVert d_{N}\right\rVert^{2}\leq\frac{1}{\gamma}\left\lVert Ld_{N}\right\rVert^{2}.$ (31)
Therefore, from (27), (30), (31), and Assumption 2,
$\left\lVert d_{k}\right\rVert^{2}$ $=\left\lVert d_{N}+d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+2d_{N}^{T}d_{R}+\left\lVert d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}$ (32)
$\leq\frac{1}{\gamma}\left\lVert Ld_{N}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}\leq\frac{1}{\gamma}\left\lVert Ld_{R}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}$
$\leq\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}\right\rVert^{2}=\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}-J_{*}^{T}F_{*}\right\rVert^{2}$
$\leq\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}L_{3}^{2}\textup{dist}^{2}(x_{k},X^{*}).$
Since $J(x)$ is continuous, it is clear that for a sufficiently small $\delta>0$, condition (13) is satisfied for all $x_{k}\in B(x^{*},\delta)$. Finally, we have
$\left\lVert d_{k}\right\rVert\leq\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{q}^{*}}L_{3}\textup{dist}(x_{k},X^{*}),$
where we recall that $\lambda_{q}^{*}$ is the smallest positive eigenvalue of $J_{*}^{T}J_{*}$. Then define $c_{1}=\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{q}^{*}}L_{3}$ and we have the result. ∎
Lemma 6.
Suppose that Assumptions 13 are valid in $B(x^{*},\delta)$, for some $\delta>0$. If $x_{k}\in B(x^{*},\delta)$, $rank(J_{k})=\ell\geq rank(J_{*})=q\geq 1$, and Assumption 4 is satisfied, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r},$ with $r\in(0,1]$, then there exists $c_{1}>0$ such that $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$.
Proof.
Consider the GSVD for the pair $(J_{k},L)$. In this case where $rank(J_{k})=\ell,rank(J_{*})=q\geq 1$ and $rank(L)=p$ with $n-q\leq p$ and $n-\ell\leq p$ (see Remark 1), we remark that:
  • For $1\leq i<n-\ell$, we have $\sigma_{i,k}$=0 and $\mu_{i,k}=1.$ This occurs because of the fact that $rank(J_{k})=\ell$ and the relationship between the singular values of the GSVD (see Theorem 1). Thus, in this case, $(\Gamma_{k})_{ii}=\frac{1}{\lambda_{k}}.$
  • For $n-\ell\leq i<n-q$, if $\sigma_{i,k}\rightarrow 0$, then $\mu_{i,k}\rightarrow 1$, and we have $(\Gamma_{k})_{ii}\rightarrow\frac{1}{\lambda_{k}}$.
  • For $n-q\leq i\leq p$, then $(\Gamma_{k})_{ii}=\frac{1}{\sigma_{i,k}+\lambda_{k}\mu_{i,k}}<\frac{1}{\sigma_{i,k}}\leq\frac{1}{\sigma^{*}_{\text{min}}},$ where $\sigma^{*}_{\text{min}}$ is the smallest positive singular value of $J(x^{*}).$
Note that, for $x_{k}\in B(x^{*},\delta)$ and due to the continuity of $\mu_{i,k}$, in the case where $n-\ell\leq i<n-q$ we can assume that $\mu_{i,k}>\frac{1}{2}$ and thus, for $\delta$ sufficiently small, we have $(\Gamma_{k})_{ii}\leq\frac{1}{4\lambda_{k}}$. Therefore,
$\left\lVert\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]\right\rVert\leq\max{\Big{[}\frac{1}{\lambda_{k}},\frac{1}{4\lambda_{k}},\frac{1}{\sigma^{*}_{\text{min}}},1\Big{]}}\leq\frac{1}{\lambda_{k}},$ (33)
as $x_{k}\in B(x^{*},\delta)$ and $\left\lVert J_{k}^{T}F_{k}\right\rVert<1.$
From (19) and (33), we obtain
$\left\lVert d_{k}\right\rVert$ $=\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}{J_{k}}^{T}F_{k}\right\rVert$ (34)
$=\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}\Big{(}{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})\Big{)}\right\rVert$
$\leq\left\lVert X_{k}\right\rVert^{2}\frac{1}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$
$+\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert.$
Observe from GSVD that
$X_{k}J_{k}^{T}J_{k}$ $=X_{k}X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}&0\\ 0&I_{n-p}\end{array}\right]U_{k}^{T}U_{k}\left[\begin{array}[]{cc}\Sigma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}=\left[\begin{array}[]{cc}\Sigma_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1},$
then, as $\Gamma_{k}:=(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}$,
$\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$ $\leq\left\lVert X_{k}\right\rVert\zeta\left\lVert X_{k}^{-1}\right\rVert\left\lVert x_{k}-\bar{x}_{k}\right\rVert$ (35)
$\leq\left\lVert X_{k}\right\rVert\left\lVert X_{k}^{-1}\right\rVert\left\lVert x_{k}-\bar{x}_{k}\right\rVert,$
where this last inequality comes from the fact that
$\left\lVert\left[\begin{array}[]{cc}(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}\Sigma_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]\right\rVert\leq\max\Big{\{}\max_{1\leq i\leq p}{\Big{\{}\frac{\sigma_{i,k}^{2}}{\sigma_{i,k}^{2}+\lambda_{k}\mu_{i,k}^{2}}}\Big{\}},1\Big{\}}=:\zeta\leq 1$
with $\sigma_{i,k}$ and $\mu_{i,k}$ as in (17).
Therefore, using (19), (35), (9), (20) and Lemma 3 in (34), we have
$\left\lVert d_{k}\right\rVert$ $\leq\left\lVert X_{k}\right\rVert^{2}\frac{1}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$ (36)
$+\left\lVert X_{k}\right\rVert\left\lVert X_{k}^{-1}\right\rVert\left\lVert x_{k}-\bar{x}_{k}\right\rVert$
$\leq\frac{1}{\gamma\lambda_{k}}L_{4}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{2}+\frac{1}{\gamma\lambda_{k}}\left\lVert J_{k}^{T}F(\bar{x}_{k})\right\rVert+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
Finally, from Assumption 4 and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, the result follows. ∎
4 Local Convergence
Lemmas 5 and 6 showed that $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, under specific assumptions. As we shall see in this section, such inequality is key for the local convergence analysis under an error bound condition (Assumption 3).
Apart from the next lemma which is an intermediate result, the remaining of this section is organized in subsections according to whether the Jacobian rank near the solution set is constant or not.
Throughout this section, we consider the LMMSS iteration (2)–(3) with $\alpha_{k}=1$ for every $k$, that is, the “pure” LMMSS.
Lemma 7.
Let $\{x_{k}\}$ be the sequence generated by the LMMSS (with $\alpha_{k}=1$) and suppose Assumptions 1-3 hold. If $x_{k+1},x_{k}\in B(x^{*},\delta)$ and $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, then
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||$ (37)
$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||,$
where $L_{5}:=L_{0}L_{2}(1+c_{1})(2+c_{1})$.
Proof.
For any $x,y\in\mathbb{R}^{n}$, the reverse triangle inequality implies
$\|\nabla\phi(y)-\nabla\phi(x)-J(x)^{T}J(x)(y-x)\|\geq\|\nabla\phi(y)\|-\|\nabla\phi(x)-J(x)^{T}J(x)(y-x)\|.$
Thus, for $y=x_{k+1}$, $x=x_{k}$, $x_{k+1},x_{k}\in B(x^{*},\delta)$ and using Lemma 1, we have:
$\|\nabla\phi(x_{k+1})\|$ $\leq\|\nabla\phi(x_{k+1})-\nabla\phi(x_{k})-J_{k}^{T}J_{k}d_{k}\|+\|\nabla\phi(x_{k})+J_{k}^{T}J_{k}d_{k}\|$ (38)
$\leq L_{4}\|d_{k}\|^{2}+\|(J_{k}-J_{k+1})^{T}F_{k+1}\|+\|J_{k}^{T}F_{k}+J_{k}^{T}J_{k}d_{k}\|.$
From the error bound condition (Assumption 3), the definition of the LMMSS iteration and (38), we get
$\omega\textup{dist}(x_{k+1},X^{*})\leq L_{4}\|d_{k}\|^{2}+\|(J_{k}-J_{k+1})^{T}F_{k+1}\|+\lambda_{k}\left\lVert L\right\rVert^{2}\left\lVert d_{k}\right\rVert.$ (39)
Now, rewriting (8), we have
$\|(J_{k}-J_{k+1})^{T}F_{k+1}\|$ $\leq L_{0}L_{2}\|x_{k}-\bar{x}_{k}\|\|x_{k+1}-\bar{x}_{k}\|+\|J_{k}^{T}F(\bar{x}_{k})\|$ (40)
$+L_{0}L_{2}\|x_{k+1}-\bar{x}_{k}\|^{2}+\|J_{k+1}^{T}F(\bar{x}_{k})\|.$
Furthermore, since $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, it holds
$\|x_{k+1}-\bar{x}_{k}\|=\|x_{k}-\bar{x}_{k}+x_{k+1}-x_{k}\|\leq\|x_{k}-\bar{x}_{k}\|+\|d_{k}\|\leq(1+c_{1})\|x_{k}-\bar{x}_{k}\|.$ (41)
Thus, using (40) and (41) in (39) yields
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq L_{4}\|d_{k}\|^{2}+L_{0}L_{2}\|x_{k}-\bar{x}_{k}\|\|x_{k+1}-\bar{x}_{k}\|+\|J_{k}^{T}F(\bar{x}_{k})\|$ (42)
$+L_{0}L_{2}\|x_{k+1}-\bar{x}_{k}\|^{2}+\|J_{k+1}F(\bar{x}_{k})\|+\lambda_{k}\left\lVert L\right\rVert^{2}\left\lVert d_{k}\right\rVert$
$\leq L_{4}\|d_{k}\|^{2}+L_{0}L_{2}(1+c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+\|J_{k}^{T}F(\bar{x}_{k})\|$
$+L_{0}L_{2}(1+c_{1})^{2}\|x_{k}-\bar{x}_{k}\|^{2}+\|J_{k+1}^{T}F(\bar{x}_{k})\|+\lambda_{k}\left\lVert L\right\rVert^{2}\left\lVert d_{k}\right\rVert$
$\leq L_{4}c_{1}^{2}\|x_{k}-\bar{x}_{k}\|^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|$
$+L_{0}L_{2}(1+c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+L_{0}L_{2}(1+c_{1})^{2}\|x_{k}-\bar{x}_{k}\|^{2}$
$+\|J_{k}^{T}F(\bar{x}_{k})\|+\|J_{k+1}^{T}F(\bar{x}_{k})\|.$
4.1 Constant Rank
In this section, we assume that $1\leq\text{rank}(J(x^{*}))=q\leq\min\{n,m\}$, and $\text{rank}(J(x))=\text{rank}(J(x^{*}))$ for each $x\in B(x^{*},\delta)$. From Lemma 5, we recall that $\left\|d_{k}\right\|\leq c_{1}\text{dist}(x_{k},X^{*})$, with $c_{1}=\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{p}^{*}}L_{3}$, provided that $\lambda_{k}>0,$ for $x_{k}\neq x^{*}$.
Next, we present two lemmas that will aid in proving the convergence theorem under Assumption 4.
Lemma 8.
Suppose that Assumptions 1-3 are valid, and $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$. Additionally, assume that Assumption 4 is verified, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, with
$\varepsilon\leq\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r},$ (43)
with $\eta\in(0,1)$ and $L_{6}=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$, then
$\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
Proof.
We know that $J(x)$ has constant rank in $B(x^{*},\delta)$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, and Assumptions 4 is satisfied. Thus, from (37) and (7), we have
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||$ (44)
$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||$
$\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{2}$
$+C\|x_{k}-\bar{x}_{k}\|^{1+r}+C\|x_{k+1}-\bar{x}_{k}\|^{1+r}$
$\leq(L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+C\|x_{k}-\bar{x}_{k}\|^{1+r}$
$+C(1+c_{1})^{1+r}\|x_{k}-\bar{x}_{k}\|^{1+r}$
$=L_{6}\|x_{k}-\bar{x}_{k}\|^{2}+(1+(1+c_{1})^{1+r})C\|x_{k}-\bar{x}_{k}\|^{1+r}$
$\leq L_{6}\|x_{k}-\bar{x}_{k}\|^{1+r}+(1+(1+c_{1})^{1+r})C\|x_{k}-\bar{x}_{k}\|^{1+r}$
$=\Big{[}L_{6}+(1+(1+c_{1})^{1+r})C\Big{]}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{1+r}$
$=:\tilde{C}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{1+r}=\tilde{C}\textup{dist}(x_{k},X^{*})^{1+r}$
where $L_{6}:=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$ and $\tilde{C}:=L_{6}+(1+(1+c_{1})^{1+r})C$.
From (44), we have
$\textup{dist}(x_{k+1},X^{*})\leq\dfrac{\tilde{C}}{\omega}\textup{dist}(x_{k},X^{*})^{1+r}\leq\dfrac{\tilde{C}}{\omega}\varepsilon^{r}\textup{dist}(x_{k},X^{*})$
and, given $\eta\in]0,1[$, for $\varepsilon\leq\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r}$ we obtain
$\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*}).$
Lemma 9.
Suppose that the assumptions of Lemma 8 are satisfied, and
$\varepsilon=\min\left\{\frac{\frac{\delta}{2}}{1+\frac{c_{1}}{1-\eta}},\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r}\right\}.$ (45)
If $x_{0}\in B(x^{*},\varepsilon)$, then $x_{k+1}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{k},X^{*})\leq\varepsilon$, for every $k\in\mathbb{N}$.
Proof.
The proof is done by induction. For $k=0$, we have $\textup{dist}(x_{0},X^{*})\leq\|x_{0}-x^{*}\|\leq\varepsilon$ and
$\|x_{1}-x^{*}\|\leq\|x_{1}-x_{0}\|+\|x_{0}-x^{*}\|\leq\|d_{0}\|+\varepsilon\leq c_{1}\textup{dist}(x_{0},X^{*})+\varepsilon\leq(1+c_{1})\varepsilon<\frac{\delta}{2}.$
For $k\geq 1$, assume that $x_{i}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{i-1},X^{*})\leq\varepsilon$ for $i=1,...,k$. Then, from Lemma 8 applied to $x_{k-1}$ and $x_{k}$, we obtain
$\|x_{k}-\bar{x}_{k}\|=\textup{dist}(x_{k},X^{*})\leq\eta\textup{dist}(x_{k-1},X^{*})\leq\eta\varepsilon<\varepsilon.$
Moreover, by Lemma 8 and the induction hypothesis for $i\leq k$, we have
$\textup{dist}(x_{i},X^{*})\leq\eta\textup{dist}(x_{i-1},X^{*})\leq\cdots\leq\eta^{i}\textup{dist}(x_{0},X^{*})\leq\eta^{i}\varepsilon<\varepsilon.$
Thus, using Lemma 5,
$\|x_{k+1}-x^{*}\|$ $\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}\|d_{i}\|\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}c_{1}\textup{dist}(x_{i},X^{*})$ (46)
$\leq(1+c_{1})\varepsilon+c_{1}\varepsilon\sum_{i=1}^{\infty}\eta^{i}=\Big{(}1+\frac{c_{1}}{1-\eta}\Big{)}\varepsilon\leq\frac{\delta}{2}.$
Theorem 2.
Suppose that Assumptions 1-3 are valid, and $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$. Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ for all $k$. If Assumption 4 is verified and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (45), then $\{\textup{dist}(x_{k},X^{*})\}$ converges superlinearly to zero. Moreover, the sequence $\{x_{k}\}$ converges to a solution $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
Proof.
From (44) we obtain
$\dfrac{\textup{dist}(x_{k+1},X^{*})}{\textup{dist}(x_{k},X^{*})}\leq\dfrac{\tilde{C}}{\omega}\textup{dist}(x_{k},X^{*})^{r}$
and since $\textup{dist}(x_{k},X^{*})^{r}\rightarrow 0$ (due to Lemmas 8 and 9) we have that $\{\textup{dist}(x_{k},X^{*})\}$ converges to zero superlinearly.
Once $\textup{dist}(x_{k},X^{*})$ converges to zero and $x_{k}\in B(x^{*},\frac{\delta}{2})$ for every $k$, it remains to show that $\{x_{k}\}$ converges. From Lemma 5 and Lemma 8, we have
$\|d_{k}\|\leq c_{1}\text{dist}(x_{k},X^{*})\leq c_{1}\eta^{k}\text{dist}(x_{0},X^{*})\leq c_{1}\varepsilon\eta^{k},$
for every $k\geq 1$. Thus, for any positive integers $\ell$ and $q$, with $\ell\geq q$,
$\|x_{\ell}-x_{q}\|\leq\sum_{i=q}^{\ell-1}\|d_{i}\|\leq\sum_{i=q}^{\infty}\|d_{i}\|\leq c_{1}\varepsilon\sum_{i=q}^{\infty}\eta^{i},$
implying that $\{x_{k}\}\subset\mathbb{R}^{n}$ is a Cauchy sequence and, therefore, converges. ∎∎
Remark 3.
In particular, when $r=1$ in Assumption 4, $\textup{dist}(x_{k},X^{*})$ converges to zero quadratically.
4.2 Diminishing Rank
Since the matrix rank is a lower semicontinuous function, for a given sequence $\{x_{k}\}$ converging to $x^{*}\in X^{*}$ it may happen that $\text{rank}J(x_{k})>\text{rank}J(x^{*})$, for every $k$ sufficiently large. We refer to this situation as “diminishing rank” scenario. In this case the convergence results will depend on the specific choice $\lambda_{k}=\|J(x_{k})^{T}F(x_{k})\|^{r^{\prime}}$, with $r^{\prime}\leq r$ from Assumption 4. To simplify the exposition, we consider $r^{\prime}=r$ in the results that follow.
Lemma 10.
Suppose that Assumptions 1-3 and 4 are valid, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
$\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r}$ (47)
with $\eta\in(0,1)$, then
$\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
Proof.
Note that from Assumption 4 and Lemma 6, for $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, we have $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$, with
$c_{1}=\frac{1}{\gamma\omega^{r}}(L_{4}+C)+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}.$ (48)
From Lemma 7, Assumption 4, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, (10), $\textup{dist}(x_{k},X^{*})\leq\delta/2<1/2$ and assuming that $\textup{dist}(x_{k},X^{*})<\varepsilon$, we have
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||+$ (49)
$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||$
$\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{1+r}$
$+C\|x_{k}-\bar{x}_{k}\|^{1+r}+C\|x_{k+1}-\bar{x}_{k}\|^{1+r}$
$\leq(L_{4}c_{1}^{2}+L_{5}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1})\|x_{k}-\bar{x}_{k}\|^{2}$
$+(1+(1+c_{1})^{1+r})C\|x_{k}-\bar{x}_{k}\|^{1+r}$
$=[L_{7}+(1+(1+c_{1})^{1+r})C]\|x_{k}-\bar{x}_{k}\|^{1+r}$
$\leq[L_{7}+(1+(1+c_{1})^{1+r})C]\varepsilon^{r}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$
$:=\hat{C}\varepsilon^{r}\left\lVert x_{k}-\bar{x}_{k}\right\rVert,$
with $c_{1}$ as in (48) and $L_{7}=L_{4}c_{1}^{2}+L_{5}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}$.
Therefore, for $\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r},$ we have $\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*})$. ∎∎
Using an appropriate $\varepsilon>0$, and Lemma 10 , the proofs of the next results are analogous to those of Lemma 9 and Theorem 2, respectively.
Lemma 11.
Suppose that the Assumptions of Lemma 10 are satisfied, and consider $\varepsilon$ as
$\varepsilon=\min\left\{\frac{\frac{\delta}{2}}{1+\frac{c_{1}}{1-\eta}},\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r}\right\}.$ (50)
If $x_{0}\in B(x^{*},\varepsilon)$, then $x_{k+1}\in B(x^{*},\frac{\delta}{2})$, and $\textup{dist}(x_{k},X^{*})\leq\varepsilon$ for every $k\in\mathbb{N}$.
Theorem 3.
Suppose that Assumptions 1-3 and 4 are valid. Let $\{x_{k}\}$ be generated by LMMSS with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, for all $k$. If $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ as in (50), then $\{dist(x_{k},X^{*})\}$ converges superlinearly to zero. Furthermore, the sequence $\{x_{k}\}$ converges to the solution $\bar{x}\in X^{*}\cap B(x^{*},\delta/2).$
Remark 4.
The main difference from Theorem 2 is that when Assumption 4 holds with $10<r<$ 1 we cannot choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ anymore; $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ is essential for the convergence proof in this case. This becomes clear from the proof of Lemma 6; see inequality (36). In fact, we could choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r^{\prime}}$ with $r^{\prime}\leq r$ that all analysis remains valid with minor modifications.
5 Global Convergence
Now we consider Algorithm 1, a version of LMMSS globalized with line-search for non-zero residue nonlinear least-squares. It is worth to point out that Algorithm 1 differs from the algorithm proposed in [5] in the choice of the LM parameter $\lambda_{k}$ as well as the full step ($\alpha_{k}=1$) acceptance criterion.
Input: $\nu,\zeta,\vartheta,\theta\in(0,1)$, $M>0$ $F$, $J$, $L$, and $x_{0}\in\Omega$.
Set $\lambda_{0}=\left\lVert\nabla\phi({x}_{0})\right\rVert$, $k=0$.
  1. 1.
    If $\|\nabla\phi(x_{k})\|=0$, stop and return $\bar{x}=x_{k}$.
  2. 2.
    Compute $d_{k}$, a solution of $(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d=-J_{k}^{T}F_{k}$.
  3. 3.
    If $\left\lVert\nabla\phi(x_{k}+d_{k})\right\rVert\leq\vartheta\left\lVert\nabla\phi(x_{k})\right\rVert$, set $\alpha_{k}=1$ and go to Step 6.
  4. 4.
    If $\|d_{k}\|>M$ or $-\nabla\phi(x_{k})^{T}d_{k}<\theta\|\nabla\phi(x_{k})\|^{2}$, set
    $d_{k}=-(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}J_{k}^{T}F_{k}.$
  5. 5.
    Choose $m$ as the smallest non-negative integer such that
    $\phi(x_{k}+\zeta^{m}{d}_{k})-\phi({x}_{k})\leq\nu\zeta^{m}\nabla\phi({x}_{k})^{T}{d}_{k}$
    Set $\alpha_{k}=\zeta^{m}$.
  6. 6.
    Update ${x}_{k+1}={x}_{k}+\alpha_{k}{d}_{k}$, $\lambda_{k+1}=\left\lVert\nabla\phi(x_{k+1})\right\rVert$, $k\leftarrow k+1$,
and go to Step 1.
Algorithm 1 LMMSS with line-search
Initially, we will demonstrate that the sequence of directions $d_{k}$ generated by Algorithm 1 is gradient-related (see [4, Sec. 1.2]). Subsequently, we will establish that any limit point of the sequence produced by this algorithm is a stationary point for (1), regardless of the initial point.
In this section, we assume that Assumption 4 is satisfied with $r=1$ to simplify notation and proofs.
We recall the definition of gradient-related directions from [4, Sec. 1.2].
Definition 1 (Gradient-related).
Let $\{x_{k}\}$ and $\{d_{k}\}$ be sequences in $\mathbb{R}^{n}$. The sequence $\{d_{k}\}$ is said to be gradient-related to $\{x_{k}\}$ if, for each subsequence $\{x_{k}\}_{k\in\mathcal{K}}$ (with $\mathcal{K}\subseteq\mathbb{N}$) converging to a non-stationary point of $\phi$, the corresponding subsequence $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded and satisfies
$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}<0.$ (51)
Proposition 1.
Let $\{d_{k}\}$, $\{x_{k}\}$ be sequences generated by Algorithm 1 starting from $x_{0}\in\mathbb{R}^{n}$. Suppose that Assumption 1 holds in the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$. Then, $\{d_{k}\}$ is gradient-related.
Proof.
Suppose that Algorithm 1 generates an infinite sequence $\{x_{k}\}$, i.e., $\|\nabla\phi(x_{k})\|=\|J_{k}^{T}F_{k}\|>0$, for each $k$, then, consequently, $\left\lVert F(x_{k})\right\rVert>0$, and let $\{x_{k}\}_{k\in\mathcal{K}}$, $\mathcal{K}\subseteq\mathbb{N}$, be a subsequence converging to $x_{\infty}$, a non-stationary point for $\phi$, i.e., $\nabla\phi(x_{\infty})\neq 0$. From now on, consider $k\in\mathcal{K}$. From (18) and the definition of $d_{k}$, it follows that
$\nabla\phi(x_{k})^{T}d_{k}$ $=-d_{k}^{T}(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}$
$=-d_{k}^{T}X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}d_{k}$
$:=-d_{k}^{T}X_{k}^{-T}\Lambda_{k}X_{k}^{-1}d_{k}.$
By Ostrowski’s inertia law, we have
$\lambda_{n}(X_{k}^{-T}\Lambda_{k}X_{k}^{-1})=\sigma_{n}^{2}(X_{k}^{-1})\lambda_{n}(\Lambda_{k}),$
where, with an abuse of notation, $\lambda_{n}(B)$ and $\sigma_{n}(B)$ represent the smallest eigenvalue and the smallest singular value of a matrix $B$, respectively. Then, from Lemma 3 it follows that
$\lambda_{n}(X_{k}^{-T}\Lambda_{k}X_{k}^{-1})\geq\gamma\lambda_{n}(\Lambda_{k})\geq\gamma\min\{1,\lambda_{k}\},$
where, in the last inequality, we use the fact that $\sigma_{i,k}^{2}+\mu_{i,k}^{2}=1$. Since $x_{\infty}$ is non-stationary, and $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert$, there exists a $\underline{\lambda}>0$ such that for $k\in{\cal K}$, $\lambda_{k}\geq\underline{\lambda}$. Thus, we obtain
$\nabla\phi(x_{k})^{T}d_{k}\leq-\gamma\min\{1,\underline{\lambda}\}\|d_{k}\|^{2}.$ (52)
On the other hand,
$\|\nabla\phi(x_{k})\|$ $=\|(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}\|\leq\|J_{k}^{T}J_{k}+\lambda_{k}L^{T}L\|\|d_{k}\|$
$\leq\left(\|J_{k}\|^{2}+\lambda_{k}\|L\|^{2}\right)\|d_{k}\|.$
By the continuity of $F(x)$ and $J(x)$ and the convergence of the subsequence $\{x_{k}\}_{\mathcal{K}}$, there exist $L_{J}>0$ and $M_{0}>0$ such that $\|J(x_{k})^{T}F(x_{k})\|\leq M_{0}$ and $\left\lVert J(x_{k})\right\rVert\leq L_{J}$ for each $k$.
Thus, since $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert$, we have
$\|\nabla\phi(x_{k})\|\leq\left(L_{J}^{2}+M_{0}\|L\|^{2}\right)\|d_{k}\|=:M_{1}\|d_{k}\|,\quad\forall k\in{\cal K}.$ (53)
From (53) and (52), we obtain
$\nabla\phi(x_{k})^{T}d_{k}\leq-\frac{\gamma\min\{1,\underline{\lambda}\}}{M_{1}^{2}}\|\nabla\phi(x_{k})\|^{2},\quad\forall k\in{\cal K},$
which implies that
$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}\leq-\frac{\gamma\min\{1,\underline{\lambda}\}}{M_{1}^{2}}\|\nabla\phi(x_{\infty})\|^{2}<0.$
Now we just need to show that $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded. Recall from (19) that
$\|d_{k}\|\leq\|X_{k}\|^{2}\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|,\quad\forall k\geq 0,$ (54)
with $\Gamma_{k}=(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}$.
From Lemma 3, we have $\|X_{k}\|\leq 1/\sqrt{\gamma}$. Now let us analyze the term $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|$. From Lemma 4, it follows that
$\|\Gamma_{k}\|\leq\dfrac{1}{\lambda_{k}},\quad\text{if}\quad 0<\lambda_{k}<1,\quad\text{and}\quad\|\Gamma_{k}\|\leq 1,\quad\text{if}\quad\lambda_{k}\geq 1.$ (55)
We separate into two cases to analyze such upper bound.
  • (a)
    If $10<\lambda_{k}<$ 1, then $\max\{\|\Gamma_{k}\|,1\}\leq\dfrac{1}{\lambda_{k}}$. Since $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, we have
    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\dfrac{1}{\lambda_{k}}\|J_{k}^{T}F_{k}\|=\dfrac{1}{\|J_{k}^{T}F_{k}\|}\|J_{k}^{T}F_{k}\|=1.$
  • (b)
    If $\lambda_{k}\geq 1$, then $\max\{\|\Gamma_{k}\|,1\}\leq 1$. Thus, $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq M_{0}$.
Therefore, from (a) and (b), we conclude that
$\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\max\{1,M_{0}\},\quad\forall k\in{\cal K}.$
and
$\|d_{k}\|\leq\frac{1}{\gamma}\max\{1,M_{0}\}=:M_{2},\quad\forall k\in{\cal K}.$ (56)
Inequalities (52) and (56) show that for $k\in{\cal K}$, $d_{k}$ is the one determined in Step 2 of Algorithm 1 because it passes the seize of Step 4 for $M=M_{2}$ and $\theta=\gamma\min\{1,\underline{\lambda}\}/M_{1}^{2}$. Thus, the proof is complete. ∎∎
Remark 5.
The above proposition shows that LMMSS directions generated by Algorithm 1 are gradient-related as long as the completeness condition (Assumption 1) holds at every iterate – which also ensures that the solution of (2) is unique. Nevertheless, apart from special cases, e.g when the Jacobian $J(x)$ is full-rank everywhere, we cannot ensure that Assumption 1 holds at every $x_{k}$ generated by Algorithm 1. Hence, we safeguard the algorithm by using the classic LMM direction whenever the LMMSS direction fails to fulfill sufficient conditions for gradient-related property, namely, $\|d_{k}\|\leq M$, for $M>0$ and $-\nabla\phi(x_{k})^{T}d_{k}\geq\theta\|\nabla\phi(x_{k})\|^{p_{1}}$, for $\theta\in(0,1)$ and $p_{1}>0$.
If such conditions are not satisfied, we take $d_{k}=-(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}J_{k}^{T}F_{k}$ which yields gradient-related directions. In fact, from
$(J_{k}^{T}J_{k}+\lambda_{k}I)d_{k}=-\nabla\phi(x_{k}),$
we obtain $-\nabla\phi(x_{k})^{T}d_{k}=\|J_{k}d_{k}\|^{2}+\lambda_{k}\|d_{k}\|^{2}\geq\lambda_{k}\|d_{k}\|^{2},$ and since
$\|\nabla\phi(x_{k})\|=\|(J_{k}^{T}J_{k}+\lambda_{k}I)d_{k}\|\leq\|J_{k}^{T}J_{k}+\lambda_{k}I\|\|d_{k}\|\leq M_{3}\|d_{k}\|,$
where $M_{3}=L_{J}^{2}+M_{0}$. We conclude that
$-\nabla\phi(x_{k})^{T}d_{k}\geq\lambda_{k}\|d_{k}\|^{2}\geq\frac{\lambda_{k}}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2}=\frac{1}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{3}$
if we take $\lambda_{k}=\|\nabla\phi(x_{k})\|$. Moreover,
$\|d_{k}\|=\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\nabla\phi(x_{k})\|$ $\leq\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\|\|\nabla\phi(x_{k})\|\leq 1,$
which shows that our “safeguard directions” are also gradient-related.
Now that it is proven that Algorithm 1 generates a sequence of directions $\{d_{k}\}$ that is gradient-related, using [4, Proposition 1.2.1], global convergence can be established.
Theorem 4.
Let $\{x_{k}\}$ be a sequence generated by Algorithm 1. Then, every limit point $\hat{x}$ of $\{x_{k}\}$ is such that $\nabla\phi(\hat{x})=0$.
Proof.
Let $K_{1}=\{k\in\mathbb{N}\mid\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|\leq\vartheta\|J(x_{k})^{T}F(x_{k})\|\}$. If $K_{1}$ is infinite, it follows that $\|J(x_{k})^{T}F(x_{k})\|\rightarrow 0$, and therefore any limit point $\hat{x}$ of $\{x_{k}\}$ is such that $J(\hat{x})^{T}F(\hat{x})=0$, hence $\nabla\phi(\hat{x})=0$. Otherwise, if $K_{1}$ is finite, let us assume, without loss of generality, that $\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|>\vartheta\|J(x_{k})^{T}F(x_{k})\|$, for each $k$, such that the step size is chosen to satisfy the Armijo condition. Since, by Proposition 1 and Remark 5, the directions of Algorithm 1 are gradient-related, it follows from [4, Proposition 1.2.1] that any limit point $\hat{x}$ of $\{x_{k}\}$ is a stationary point of $\phi(x)$. ∎∎
The next result establishes the connection between global and local convergence by showing that, under certain conditions, $\alpha_{k}=1$ for all $k$ sufficiently large and then Algorithm 1 becomes the “pure” LMMSS and the results of Section 4 are applicable.
Theorem 5.
Let $\{x_{k}\}$ be generated by Algorithm 1, using $x_{0}\in\mathbb{R}^{n}$ as the initial point. Suppose that Assumptions 1-3 and Assumption 4, with $r=1$, hold at the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$. Moreover, assume that ${\cal L}_{0}$ is compact. Then, for every $k$ sufficiently large $\alpha_{k}=1$ and the sequence $\{\textup{dist}(x_{k},X^{*})\}$ converges to 0 quadratically.
Proof.
Note that by the definition of Algorithm 1 $\{\phi(x_{k})\}$ is decreasing. Starting from $x_{0}$, we conclude that the sequence $\{x_{k}\}$ remains in ${\cal L}_{0}$, a compact set. Therefore, there exists a convergent subsequence, as per Theorem 4, where every limit point of this subsequence is stationary. By removing this convergent subsequence, we can repeat this process successively for the remaining sequence, which remains in a compact set, thereby concluding that $\{x_{k}\}$ is the disjoint and enumerable union of convergent subsequences, each of which has a stationary point as limit.
Thus, for every $k$ sufficiently large, the norm of the gradient at $x_{k}$ will be sufficiently small. In other words, there is a positive integer $k_{0}$ such that
$\|J(x_{k})^{T}F(x_{k})\|\leq\frac{\omega^{3}\vartheta}{\tilde{C}L_{3}},\quad\text{for all}\quad k\geq k_{0}.$ (57)
For $k\geq k_{0}$, let $\bar{x}_{k+1}\in X^{*}$ such that $\|x_{k+1}-\bar{x}_{k+1}\|=\textup{dist}(x_{k+1},X^{*})$. Then, from (7), Assumption 3,(44) and (57), it follows that
$\|J(x_{k+1})^{T}F(x_{k+1})\|$ $=\|J(x_{k+1})^{T}F(x_{k+1})-J(\bar{x}_{k+1})^{T}F(\bar{x}_{k+1})\|$
$\leq L_{3}\|x_{k+1}-\bar{x}_{k+1}\|$
$=L_{3}\textup{dist}(x_{k+1},X^{*})\leq\frac{\tilde{C}}{\omega}L_{3}\textup{dist}(x_{k},X^{*})^{2}$
$\leq\frac{\tilde{C}L_{3}\|J(x_{k})^{T}F(x_{k})\|}{\omega^{3}}\|J(x_{k})^{T}F(x_{k})\|$ (58)
$\leq\vartheta\|J(x_{k})^{T}F(x_{k})\|,$
proving that $\alpha_{k}=1$ for each $k\geq k_{0}$. To obtain the quadratic convergence of $\{\textup{dist}(x_{k},X^{*})\}_{k\geq k_{0}}$ to 0, one may apply Theorem 3 when we have a diminishing rank or from Theorem 2, when $rank(J(x))=rank(J(x^{*}))\geq 1$. ∎∎
Remark 6.
We observe that in the case of zero residue a previous result [13, Theorem 3.1] assumed that the limit point $x^{*}$ is such that $F(x^{*})=0$. Also, in the case of unconstrained optimization, in order to prove a result similar to Theorem 5, the required assumption is that the Hessian at $x^{*}$ is positive definite; however, such assumption would imply that $x^{*}$ is an isolated stationary point. As we want to address the case of non-zero residue and possibly non-isolated stationary points, we considered the boundness of the level set instead of these other two conditions.
Remark 7.
In the scenario of diminishing rank, if Assumption 4 is satisfied with $10<r<$ 1, it is necessary to adjust the parameter $\lambda_{k}$ of Algorithm 1 to $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$. Furthermore, when this hypothesis holds, the above results follow with minor changes in some constants, and we can show there exists $k_{0}$ such that $\alpha_{k}=1,\forall k\geq k_{0}$ and that $\textup{dist}(x_{k},X^{*})$ goes to zero superlinearly.
6 Illustrative examples
The purpose of this section is to illustrate our convergence results by presenting some examples of non-zero residue nonlinear least-squares problems. In this section we use the shorthand notation (A1) – (A4) for Assumptions 14, respectively.
Example 1. Let us start with an example from [2, Example 5.1] with non-isolated minimizers (stationary-points). Consider the residual function
$F(x)=(x_{1}^{2}+x_{2}^{2}-1,x_{1}^{2}+x_{2}^{2}-9)^{T}.$
The set of minimizers (stationary-points) $X^{*}$ consists of points satisfying $x_{1}^{2}+x_{2}^{2}-5=0$, thus $\text{dist}(x,X^{*})=|\|x\|-\sqrt{5}|$. Moreover, for any $x^{*}\in X^{*}$, $\phi(x^{*})=\frac{1}{2}\|F(x^{*})\|^{2}=\frac{1}{2}(4^{2}+(-4)^{2})=16$.
It is straightforward to observe that
$J(x)=\left[\begin{array}[]{cc}2x_{1}&2x_{2}\\ 2x_{1}&2x_{2}\end{array}\right]$
has rank $1$ everywhere except for the origin. Moreover, as discussed in [2], Assumptions (A2), (A3) and (A4) with $r=1$ hold as well.
By considering LMMSS with the scaling matrix $L=[-1\quad 1]$, Assumption (A1) holds everywhere except for the line $x_{2}=-x_{1}$.
Hence, we are under conditions to apply the results of Section 4.1. As expected from Theorem 2, in Table 1, we can observe the local quadratic convergence of $\text{dist}(x_{k},X^{*})$ to zero starting from two distinct starting points in a neighborhood of $x^{*}=(0,\sqrt{5})^{T}\in X^{*}$.
Table 1: Stopping criterion $\|J_{k}^{T}F_{k}\|<10^{-8}$ with $x_{0}=(0,\sqrt{5}+0.03)^{T}$ as starting point (left), and $x_{0}=(0.01,\sqrt{5}-0.01)^{T}$ (right).
$k$ dist($x_{k},X^{*}$) $\left\lVert J_{k}^{T}F_{k}\right\rVert$ $k$ dist($x_{k},X^{*}$) $\left\lVert J_{k}^{T}F_{k}\right\rVert$
0 $1.3506\times 10^{-1}$ $1.2242$ 0 $4.4521\times 10^{-2}$ $3.9643\times 10^{-1}$
1 $1.7762\times 10^{-2}$ $1.5890\times 10^{-1}$ 1 $1.9821\times 10^{-4}$ $1.7729\times 10^{-3}$
2 $3.2402\times 10^{-7}$ $2.8982\times 10^{-6}$ 2 $3.8598\times 10^{-9}$ $3.4523\times 10^{-8}$
3 $1.1546\times 10^{-14}$ $1.0659\times 10^{-13}$
This example is also useful to illustrate the effect of using $L^{T}L\neq I$ and the globalization mechanisms of Algorithm 1.
Figure 1a depicts the trajectory of LMM ($L^{T}L=I$) and the one of LMMSS (with $L=[-1\quad 1]$) both starting from $x_{0}=(2,4)^{T}$. As we can see from the figure, while LMM dives perpendicularly to the set $X^{*}$ (observe the level curves in the figure), LMMSS trajectory is parallel to the null space of $L$, and they end up at different stationary points. As exposed in [5] for parameter identification problems, an advantage of using Algorithm 1 is that in a given problem with non-isolated solutions, we can choose $L$ that enforces a specific property. In this geometric example, by using $L=[-1,1]$, we aim for the method to prioritize stationary points where $\|L(x-x_{0})\|=|x_{2}-x_{1}-2|$ is as small as possible. In [5], the authors employed $L$ as discrete versions of derivative operators in order to promote smoothness in the obtained solutions for parameter identification problems in heat conduction. Moreover, Figure 1b enforces this aspect by showing that the presence of a well-thought choice of $L$ may conduce the iterations faster and in a more accurate manner than LMM, while it stimulates a desired property.
(a) Iterations and level curves of the objective function.
(b) Values of $\|J_{k}^{T}F_{k}\|$ until the stopping criterion is achieved.
Figure 1: Plot of iterations for LMM and LMMSS for Example 1, with starting point $(2,4)$.
The last aspect we shall discuss with this example is about the completeness condition and the global convergence mechanisms of Algorithm 1. As mentioned before, (A1) holds everywhere except for the line $x_{2}=-x_{1}$. LMMSS also tries to move parallel to the null space of $L$. Figure 2a illustrates what happens if we apply LMMSS without the safeguard Step 4, using $x_{0}=(-1,3)^{T}$. The displacements $x_{k}-x_{0}$ are in ${\cal N}(L)$, but the hyperplane $x_{0}+{\cal N}(L)$ does not intersect $X^{*}$. As a result, LMMSS minimizes $\phi(x)$ in $x_{0}+{\cal N}(L)$ approximating a point $\hat{x}$ where $\nabla\phi(\hat{x})\perp{\cal N}(L)$ so that the directions $d_{k}$ become orthogonal to $-\nabla\phi(x_{k})$ loosing the descent property and the algorithm crashes not satisfying the Armijo condition. Notice that $\hat{x}$ is such that $\hat{x}_{2}=-\hat{x}_{1}$ and (A1) fails. On the other hand, we use Algorithm 1 with the safeguard step, Step 4 corrects this behavior and when $x_{k}$ approaches $\hat{x}$ the classic LMM direction is employed instead; see Figure 2b. This allows the algorithm to move on and reach $X^{*}$.
(a) LMMSS without safeguard Step 4 in Algorithm 1.
(b) LMMSS with safeguard Step 4 in Algorithm 1.
Figure 2: Plot of iterations for LMM and LMMSS for Example 1, with starting point $(-1,3)$, and level curves of the objective function.
Example 2. Now we consider an example to illustrate the “diminishing rank” scenario. The residual function is given by
$F(x)=(x_{1}^{3}-x_{1}x_{2}+1,x_{1}^{3}+x_{1}x_{2}+1),$
which leads to $\phi(x)$ having an isolated global minimizer at $(-1,0)^{T}$ and a non-isolated set of local minimizers $\{x\in\mathbb{R}^{2}\mid x_{1}=0\}$ that we shall use as $X^{*}$ and, in this case, $\textup{dist}(x,X^{*})=|x_{1}|$.
This example was also analyzed in [2] where it was verified that assumptions (A2), (A3) and (A4) with $r=1$ hold in a neighborhood of $x^{*}=(0,2)^{T}$.
For assumption (A1), we point out that the null space of
$J(x)=\left[\begin{array}[]{cc}3x_{1}^{2}-x_{2}&-x_{1}\\ 3x_{1}^{2}+x_{2}&x_{1}\end{array}\right]$
is $\{0\}$ wherever $x_{1}\neq 0$ and it is $\text{span}\{(0,1)^{T}\}$ for $x\in X^{*}\setminus\{0\}$. Hence, it is clear that, regardless the choice of $L$, (A1) holds everywhere except for $X^{*}$.
In Table 2 and Figure 3 we report the behavior of LMM and LMMSS with $L=[-1\quad 1]$ starting in the neighborhood of $x^{*}=(0,2)^{T}$. From Table 2 the quadratic convergence rate is apparent for both methods. In Figure 3 we highlight the different limit points: while the classic LMM moves perpendicularly to $X^{*}$, LMMSS moves parallel to ${\cal N}(L)$, leading in this case to a stationary point with smaller $x_{2}$ coordinate.
Figure 3: Plot of iterations for LMM and LMMSS for Example 2, with starting point $(0.8,2.1)^{T}$, and level curves of the objective function. Limit points: $(1.4833\times 10^{-14},1.9915)^{T}$ for LMM and $(-7.5482\times 10^{-16},1.3377)^{T}$ for LMMSS.
Table 2: Values of dist($x_{k},X^{*}$) until the stopping criterion $\|J_{k}^{T}F_{k}\|<10^{-10}$ is achieved for Example 2 with $x_{0}=(0.8,2.1)^{T}$.
$k$ LMM LMMSS
1 $3.7143\times 10^{-1}$ $1.5307\times 10^{-1}$
2 $6.0270\times 10^{-2}$ $1.3438\times 10^{-2}$
3 $1.0055\times 10^{-3}$ $1.7991\times 10^{-4}$
4 $2.4684\times 10^{-7}$ $3.0097\times 10^{-8}$
5 $1.4833\times 10^{-14}$ $7.5482\times 10^{-16}$
Example 3. Let us consider $F(x)=(x_{1}^{2},x_{2}^{2},x_{1}+x_{2},1),$ the residual function. The corresponding least-squares function $\phi(x)$ has a unique global minimizer at $x^{*}=(0,0)^{T}$, which is the only stationary point of $\phi(x)$, $X^{*}=\{x^{*}\}$. The rank of
$J(x)=\left[\begin{array}[]{cc}2x_{1}&0\\ 0&2x_{2}\\ 1&1\\ 0&0\end{array}\right]$
is 2 for $x\neq x^{*}$ and it is 1 at $x^{*}$. Therefore the rank diminishes as we approach $X^{*}$ and (A1) holds everywhere, regardless $L$, except at $x^{*}$. Also, since $\|J(x)-J(y)\|\leq 2\|x-y\|$, (A2) is verified.
The distinguished feature of this example is that the error bound condition (A3) does not hold for every point in a neighborhood of $x^{*}$. In fact, by considering the sequence $x_{k}=(1/k,-1/k)^{T}$ it is not hard to show that
$\lim_{k\to\infty}\frac{\|\nabla\phi(x_{k})\|}{\textup{dist}(x_{k},X^{*})}=\lim_{k\to\infty}\frac{\|\nabla\phi(x_{k})\|}{\|x_{k}\|}=0,$
and thus (A3) cannot hold.
Nevertheless, for any $x\in\text{span}\{(1,1)^{T}\}={\cal N}(J(x^{*}))^{\perp}$, from the mean value theorem, we have
$\|\nabla\phi(x)-\nabla\phi(0)\|=\|\nabla^{2}\phi(\xi)x\|\geq 2\|x\|,$
where we used the fact that $(1,1)^{T}$ is an eigenvector of $\nabla^{2}\phi(\xi)$ and 2 is a lower bound for the corresponding eigenvalue. Therefore, in this example, starting at $x_{0}\in{\cal N}(J(x^{*}))^{\perp}$ and using $L$ such that ${\cal N}(L)={\cal N}(J(x^{*}))^{\perp}$, we force the sequence $\{x_{k}\}$ generated by LMMSS to stay in a subspace where (A3) holds. Also, with this choice of $L$, (A1) is satisfied everywhere, and since we have
$\|J(x)^{T}F(0)\|=0\leq K\|x-0\|^{2},$
(A4) with $r=1$ is fulfilled and quadratic convergence is achieved, as depicted in Table 3.
Table 3: Distances of $x_{k}$ to $x^{*}$ produced by LMMSS until the stopping criterion $\|J_{k}^{T}F_{k}\|<10^{-10}$ is achieved for Example 3 with $L=[-1\quad 1]$, for different starting points.
$k$ $x_{0}^{(1)}=(3,3)^{T}$ $x_{0}^{(2)}=(-2,-2)^{T}$
1 $2.0097$ $1.2571$
2 $8.0542\times 10^{-1}$ $3.8494\times 10^{-1}$
3 $1.5845\times 10^{-1}$ $2.4840\times 10^{-2}$
4 $1.9403\times 10^{-3}$ $7.6586\times 10^{-6}$
5 $3.6524\times 10^{-9}$ $2.6331\times 10^{-16}$
6 $7.4734\times 10^{-17}$
It is worth mentioning that even though (A3) does not hold for every $x$ in a neighborhood of $x^{*}$ in this example, a weaker Hölderian error bound condition [1] does hold. Precisely $\|\nabla\phi(x)\|\geq\|x\|^{3}$ or, in other words, $\nabla\phi$ is Hölder metrically subregular of order 1/3 (see [1, Definition 1]).
7 Final remarks
We have investigated convergence issues of Levenberg-Marquardt method with singular scaling (LMMSS) [5] for nonlinear least-squares problems with non-zero residue. Through a couple of modifications in the algorithm proposed in [5, Algorithm 1], mainly, the choice of Levenberg-Marquardt parameter and the line-search, we have established local and global convergence results under mild assumptions.
It was shown that LMMSS direction is well defined as long as the nullspace of the Jacobian and the nullspace of the singular scaling matrix $L^{T}L$ have trivial intersection (Assumption 1). Then, it was demonstrated local superlinear convergence rate under a local error bound condition on the gradient of the least-squares function (Assumption 3) and an assumption on the linearization error of the gradient (Assumption 4) which represents a measure of nonlinearity and residue size. These results were established for both the constant rank and diminishing rank cases.
By using an Armijo line-search and safeguard directions, it was proved that all limit points of the sequence generated by Algorithm 1 are stationary points of the least-squares function. We also presented some examples illustrating the assumptions and convergence results.
In future works we plan to consider Hölder metric subregularity [1] instead of the local error bound used here and address local convergence when a weaker form of Assumption 4 holds, namely when $\|J(x)^{T}F(z)\|={\cal O}(\|x-z\|)$ instead of $\|J(x)^{T}F(z)\|=o(\|x-z\|)$, following the lines of the analysis presented in [2].
Apart from tackling these questions, we aim to conduct computational experiments to test the algorithm, equipped with suitable stopping criteria, in certain inverse problems to assess its numerical behavior.
Acknowledgments
This work was supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]. The work of DG was partially supported by CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico), Brazil [grant number 305213/2021-0].

### v2

@@
-Convergence analysis of Levenberg-Marquardt method with Singular Scaling for non-zero residue nonlinear least-squares problems [^1] This work is supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]
+Florianópolis, 88040900, SC, Brazil
 Abstract
-This paper addresses the convergence analysis of a variant of the Levenberg-Marquardt method (LMM) designed for nonlinear least-squares problems with non-zero residue.
-This variant, called LMM with Singular Scaling (LMMSS), allows the LMM scaling matrix to be singular, encompassing a broader class of regularizers, which has proven useful in certain applications.
-In order to establish local convergence results, a careful choice of the LMM parameter is made based on the gradient linearization error, dictated by the nonlinearity and size of the residual.
-Under completeness and local error bound assumptions we prove that the distance from an iterate to the set of stationary points goes to zero superlinearly and that the iterative sequence converges.
-Furthermore, we also study a globalized version of the method obtained using line-search and prove that any limit point of the generated sequence is stationary.
+Recently, a Levenberg-Marquardt method with Singular Scaling matrix, called LMMSS, was proposed and successfully applied in parameter estimation in heat conduction problems, where the choice of suitable singular scaling matrix resulted in better quality approximate solutions than those of the classical Levenberg-Marquardt.
+In this paper, we study convergence properties of LMMSS when applied to nonzero residual nonlinear least-squares problems.
+We show that the local convergence of the iterates depends both on the control of the gradient linearization error and on a suitable choice of the regularization parameter.
+Incidentally, we show that the rate of convergence is dictated by a measure of nonlinearity and residual size, so that if such a measure goes to zero quickly enough, the convergence can be superlinear, otherwise, in general, we show that not even linear convergence can be expected if such a measure is not small enough.
+Additionally, we propose a globalized version of the method and prove that any limit point of the generated sequence is stationary for the least-squares function.
 Some examples are provided to illustrate our theoretical results.
-Keywords. Levenberg-Marquardt, Singular Scaling Matrix, Convergence analysis, Non-zero residue
+Keywords:
+Levenberg-MarquardtSingular Scaling Matrix Convergence analysis nonzero residual
+MSC:
+49M37 65K05 90C30
 1 Introduction
-In this paper, we investigate global and local convergence properties of the Levenberg–Marquardt method applied to a nonlinear least-squares (NLS) problem:
-$\min_{x\in\mathbb{R}^{n}}\quad\frac{1}{2}\|F(x)\|^{2}:=\phi(x),$ (1)
-where $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is twice continuously differentiable. In particular, we focus on the overdetermined problem, where $m\geq n$.
-Unlike previous research, which predominantly focused on the zero residue case or in cases where the rank of the Jacobian is complete [7, 6, 10, 11, 13], here we consider NLS problems where the set of zeros of $F$ is empty and its Jacobian is allowed to be rank-deficient.
-We say that problem (1) is a non-zero-residue NLS problem when the set of zeros of $F$ is empty.
-Our interest is to identify the stationary points of $\phi$, i.e., the set $X^{*}=\{x\in\mathbb{R}^{n}\mid J(x)^{T}F(x)=0\}$, where $J(x)\in\mathbb{R}^{m\times n}$ is the Jacobian of $F$ at $x$, under the assumption that $X^{*}\neq\emptyset$.
-Throughout the text we shall use $x^{*}$ to denote an arbitrary element of $X^{*}$.
-Although works such as [2, 3] explore problems with non-zero residuals and present results of local convergence of the LM method, it should be noted that the scale matrix used in these works is the identity matrix.
-In this work, in addition to assuming non-zero residue, we consider the following iteration:
+In this study we investigate convergence properties of the Levenberg–Marquardt method with singular scaling (LMMSS) from Boos2024 when applied to the nonlinear least-squares (NLS) problem:
+$\min_{x\in\mathbb{R}^{n}}\quad\phi(x):=\frac{1}{2}\|F(x)\|^{2},$ (1)
+[…]
+Unlike previous local convergence analyses Dennis_Schnabel ; FanYuan ; Gratton2007 ; Yamashita , including Boos2024 , we assume neither zero residual at a solution of (1) nor full rank of the Jacobian at such a point.
+In applied contexts, such as data fitting, parameter estimation, experimental design, and imaging problems bellavia2018elliptical ; barz2015nonlinear ; cornelio2011regularized ; deidda2014regularized ; henn2003levenberg ; landi2017limited , to name a few, admitting a nonzero residual is essential for achieving meaningful solutions.
+The LMMSS iteration is defined as follows Boos2024 :
 $(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}$ $=-J_{k}^{T}F_{k}$ (2)
-$x_{k+1}$ $=x_{k}+\alpha_{k}d_{k},\quad\forall k\geq 0,$ (3)
-where $F_{k}:=F(x_{k})$, $J_{k}:=J(x_{k})$ is the Jacobian of $F$ at $x_{k}$, $\lambda_{k}>0$, $\alpha_{k}$ is the step size, and $L^{T}L$ referred to as scaling matrix, is allowed to be singular.
+$x_{k+1}$ $=x_{k}+\alpha_{k}d_{k},$ (3)
+where $F_{k}:=F(x_{k})\in\mathbb{R}^{m}$, $J_{k}:=J(x_{k})\in\mathbb{R}^{m\times n}$ is the Jacobian of $F$ at $x_{k}$, $\lambda_{k}>0$, $\alpha_{k}$ is the step size and $L^{T}L$, referred to as scaling matrix, is allowed to be singular.
 We refer to the iteration (2)–(3) as the Levenberg-Marquardt method with Singular Scaling (LMMSS).
-This type of iteration is motivated by applications in inverse problems, as recently seen in [5], where a parameter recovery problem using LMMSS is addressed, including a local convergence analysis for the zero residue case.
-On the other hand, convergence properties of the classic LMM in the non-zero residue scenario have received recent attention in the literature [2, 3].
-These are two aspects in which our contribution differs from the existing literature.
-Our aim is to analyze the local convergence of LMM with singular scaling matrices in the non-zero residue scenario.
-Furthermore, we shall provide a globalization for the LMMSS iteration where the step-size $\alpha_{k}$ is determined by line-search and prove that any limit point of the generated sequence is stationary for the least-squares function.
-This paper is organized as follows.
-In Section 2, we start with some assumptions that will be considered throughout the manuscript.
-In Section 3, we present technical results necessary for the upcoming sections.
-We show that LMMSS iteration is well-defined and provide the local convergence results in Section 4, which are categorized into two cases based on the behavior of the rank of the Jacobian around a stationary point (constant and diminishing rank).
-Finally, Section 5 proves that, for LMMSS with an Armijo line-search scheme, every limit point of the generated sequence is a stationary point of the sum-of-squares function, thereby establishing the global convergence of this method.
-A few examples are given in Section 6 to illustrate our theoretical results.
-Final considerations are given in Section 7.
+When $L^{T}L=I$ we retrieve the classic Levenberg-Marquardt method (LMM).
+In general, problem (1) is a nonconvex optimization problem and we will limit our attention to stationary points of $\phi$. The set of stationary points will be denoted by
+$X^{*}=\{x\in\mathbb{R}^{n}\mid\nabla\phi(x)=J(x)^{T}F(x)=0\},$
+and assume that $X^{*}\neq\emptyset$.
+We are particularly interested in the case of nonisolated stationary points, with a possible change (decrease) in the rank of the Jacobian as the generated sequence of iterates $\{x_{k}\}$ approaches the set $X^{*}$.
+Given a starting point sufficiently close to $x^{*}\in X^{*}$, we are interested in the convergence analysis of the sequence generated by the local LMMSS iteration (with $\alpha_{k}=1$).
+Our main contribution consists in establishing local convergence results for LMMSS based on an error bound condition upon $\nabla\phi$ without requiring zero residual neither full rank of the Jacobian at stationary points.
+[…]
+The seminal work Yamashita showed the local convergence of LMM for systems of nonlinear equations under an error bound condition upon the norm of the residual, without relying on the assumption of the full rank of the Jacobian.
+Assuming $\lambda_{k}=\|F(x_{k})\|^{2}$, the authors established that the distance of the iterates to the solution set converges quadratically to zero.
+Later, FanYuan improved such result and showed that quadratic convergence is still attained for $\lambda_{k}=\|F(x_{k})\|^{1+r},r\in[0,1]$.
+Furthermore, they also showed that the sequence of iterates itself converges quadratically to some point in the solution set.
+More recently, a few papers have focused on establishing local convergence of LMM for nonzero residual NLS problems (i.e $\forall x\in\mathbb{R}^{n}$, $F(x)\neq 0$).
+In behling2019local , local convergence was established using an error bound on the gradient of $\phi$ and considering a possible change in the rank of Jacobian around a stationary point.
+In Bergou2020 , local convergence of LMM in the nonzero residual case was also established, but using a different error bound condition based on the distance between the residual vector at $x$ and the one at $\bar{x}$, the stationary point closest to $x$.
+Nevertheless, all the works mentioned above considered the classic LMM where $L^{T}L=I$.
+Perhaps Boos2024 was the first work to consider local convergence analysis of LMMSS (for $L^{T}L\neq I$) under an error bound condition, but considering the zero residual case.
+The motivation to use singular matrices in the form $L^{T}L$ in LMMSS iteration (2)–(3) comes from the general form of Tikhonov regularization Tikhonov95 for linear systems $Ax=b$, i.e.,
+$x_{\lambda}=(A^{T}A+\lambda L^{T}L)^{-1}A^{T}b.$
+The use of a singular scaling matrices allows us to use seminorm regularizers which can promote specific properties (inherent of the expected solution) in the approximate solution generated by method.
+Like the $\ell_{1}$-norm is used to promote sparsity, if the aimed solution is expected to be smooth, then the use of $L$ as a discrete version of derivative operator allows us to induce such smoothness in the approximate solution.
+This approach proved to be successful in ill-posed linear inverse problems Hansen1998 ; Hansen2010 ; Buccini2017 .
+Nonlinear inverse problems can also benefit from the more flexible choice of $L$ in LMMSS.
+For instance, in Boos2024 LMMSS was successfully used to estimate the 2D perfusion coefficient in a bioheat model.
+Figure 1 shows the surface plots of the 2D perfusion coefficient reconstructions: the picture on the left plots the expected solution, the one on the middle shows the reconstruction obtained by the classic LMM and the one on the right corresponds to the reconstruction obtained by LMMSS with a specific choice of $L=\mathcal{L}_{3}$, where $\mathcal{L}_{3}$ is based on a discrete version of the third order derivative operator.
+Both methods stopped after dozens of iterations based on the discrepancy principle as stopping criterion Morozov1984 .
+For this problem, it is noteworthy that although the temperature reconstruction error was the same for both methods (almost the same value of $\phi$), LMMSS provided an approximate solution closer to the expected one.
+For more details, see (Boos2024, , Section 5.1)).
+Figure 1: The picture on the left plots the expected solution, the one on the middle shows the reconstruction obtained by the classic LMM and the one on the right corresponds to the reconstruction obtained by LMMSS with a specific choice of $L$.
+Image extracted from (Boos2024, , Figure 3).
+The above example illustrates the potential of LMMSS when a suitable $L$ is chosen.
+Observe that $L$ defines the seminorm regularizer and must be chosen according to the problem at hand, based on known properties of the expected solution.
+As we already mentioned, if the solution (of the original problem, whose discretization gave rise to the nonlinear least-squares problem) is expected to be smooth, then the use of $L$ as a discrete version of derivative operators may be a good choice.
+But other examples are available in the literature, for instance, see Buccini2017 ; pes2020 ; Pes2022 and references therein.
+As far as we know, convergence analysis of LMMSS in the nonzero residual scenario has not yet been addressed in the literature.
+This is precisely the purpose of this work: for the LMMSS iteration defined in Boos2024 we aim at establishing local convergence rates to stationary points of the least-squares function in the case of nonzero residual under an error bound condition upon $\nabla\phi$, following the lines of the analysis in behling2019local .
+Of course, by considering a possibly singular $L^{T}L$ in iteration (2)–(3) makes the analysis more challenging, for the iteration itself to be well defined depends on a condition on the null spaces of $L$ and the Jacobian of $F$.
+Additionally, we propose a globalization for the LMMSS iteration where the step-size $\alpha_{k}$ is determined by a line-search satisfying an Armijo condition and the search direction is safeguarded by the classic LMM direction.
+For the resulting algorithm, we prove that any limit point of the generated sequence is stationary.
+The paper is organized as follows.
+In Section 2, we gather some necessary assumptions to show that LMMSS iteration is well-defined and for the subsequent analysis.
+In Section 3, we present some technical results and discuss the local convergence analysis in Section 4.
+As we shall see, the convergence rate is dictated by a measure of nonlinearity and residual size.
+If such a measure goes to zero fast enough, the convergence can be superlinear but, in general, we cannot expect even linear convergence if such measure is not small enough.
+In Section 5, a globalized LMMSS algorithm is proposed and its global convergence established.
+To close this theoretical study, some examples are given in Section 6 to illustrate the local and global behavior of the algorithm in nonzero residual nonlinear least-squares problems.
+Final considerations are drawn in Section 7.
 2 Assumptions
-Recall that throughout this manuscript, we assume that the set of stationary points of $\phi$ is non-empty: $X^{*}\neq\emptyset$.
-Assumption 1 below guarantees that $d_{k}$ solution of (2) is unique.
-Assumption 1.
-The matrix $L\in\mathbb{R}^{p\times n}$ has $\text{rank}(L)=p\leq n\leq m$ and there exists $\Omega\subset\mathbb{R}^{n}$ and $\gamma>0$ such that, for every $x\in\Omega$
+Recall that throughout this manuscript, we assume that $X^{*}\neq\emptyset$.
+We shall denote a generic element of $X^{*}$ by $x^{*}$.
+Given $x$, we denote by $\bar{x}$ an element of $X^{*}$ such that $\|x-\bar{x}\|=\text{dist}(x,X^{*})$.
+Also, given $x^{*}\in X^{*}$, we may use the shorthand notation $F_{*}=F(x^{*})$ and $J_{*}=J(x^{*})$.
+The nullspace of $A$ is denoted by $\mathcal{N}(A)$ and $\mathcal{R}(A)$ denotes its range.
+Definition 1
+We say that the uniform completeness condition for problem (1) with respect to $L$ holds at $\Omega\subset\mathbb{R}^{n}$, if $L\in\mathbb{R}^{p\times n}$ has $\text{rank}(L)=p\leq n\leq m$ and there exist a constant $\gamma>0$ such that, for every $x\in\Omega$
 $\left\lVert J(x)v\right\rVert^{2}+\left\lVert Lv\right\rVert^{2}\geq\gamma\left\lVert v\right\rVert^{2},\quad\forall v\in\mathbb{R}^{n}.$ (4)
-Such assumption comes from the inverse problems literature [12, 8] where it is sometimes called completeness condition.
-It was a key assumption in [5] to study the local convergence of LMMSS in the zero residue setting.
-We remark that such assumption is equivalent to $\mathcal{N}(J(x))\cap\mathcal{N}(L)=\{{\bf 0}\},\forall x\in\Omega.$
-Next assumption is quite standard in the Gauss–Newton and Levenberg–Marquardt literature and asks the Jacobian to be locally Lipschitz.
-Assumption 2.
-For any $x^{*}\in X^{*}$, there exists a constant $\delta\in(0,1)$ and $L_{0}>0$ such that
-$\|J(x)-J(y)\|\leq L_{0}\|x-y\|,$ (5)
+It is worth noting that condition (4) is equivalent to
+$\mathcal{N}(J(x))\cap\mathcal{N}(L)=\{{\bf 0}\}.$ (5)
+Condition (5) is often referred to as a completeness condition and comes from the literature of linear inverse problems Morozov1984 ; engl , where $J(x)=A,\forall x\in\mathbb{R}^{n}$.
+Observe that it clearly holds when $J(x)$ is of full rank, regardless of the choice of $L$, as it occurs in some applications involving physical problems (Boos2024, , Section 5.1).
+In the context of LMMSS iteration, if the iterates $x_{k}$ lie in a set $\Omega\subset\mathbb{R}^{n}$ where the uniform completeness condition holds, then $d_{k}$ solution of (2) exists and is unique because (4) implies $J_{k}^{T}J_{k}+\lambda_{k}L^{T}L$ is positive definite for $\lambda_{k}>0$.
+Next assumption is standard in the Gauss–Newton and Levenberg–Marquardt literature and asks the Jacobian to be locally Lipschitz.
+Assumption 1
+For some $x^{*}\in X^{*}$, there exists a constant $\delta\in]0,1[$ and $L_{0}>0$ such that
+$\|J(x)-J(y)\|\leq L_{0}\|x-y\|,$
 for all $x,y\in B(x^{*},\delta)$.
 Assumption 1 implies that:
-$\|J(y)(x-y)-(F(x)-F(y))\|\leq L_{1}\|x-y\|^{2},\quad\forall x,y\in B(x^{*},\delta)$ (6)
+$\|J(y)(x-y)-(F(x)-F(y))\|\leq L_{1}\|x-y\|^{2},\quad\forall x,y\in B(x^{*},\delta)$
 where $L_{1}=L_{0}/2$, that is, the error in the linear approximation of $F(x)$ around $y$ is $O(\|x-y\|^{2})$, for $x$ and $y$ in a neighborhood of $x^{*}$.
@@
 $+L_{0}L_{2}\|y-z\|^{2}+\|J(y)^{T}F(z)\|.$
-Lemma 1.
-[2, Lemma 2.1] If Assumption 2 is satisfied, then there exists a constant $L_{4}>0$ such that
+Lemma 1
+(behling2019local, , Lemma 2.1) If Assumption 1 is satisfied, then there exists $\delta\in]0,1[$ and a constant $L_{4}>0$ such that
 $\left\lVert\nabla\phi(y)-\nabla\phi(x)-J(x)^{T}J(x)(y-x)\right\rVert\leq L_{4}\left\lVert x-y\right\rVert^{2}+\left\lVert(J(x)-J(y))^{T}F(y)\right\rVert,$ (8)
 for all $x,y\in B(x^{*},\delta)$.
-Next, we present the local error bound condition assumed in this work.
-We remark that it is an error bound condition with respect to the set of stationary points of $\phi$ rather than the zeroes of $F$ (which is empty in the case of non-zero residue).
-Assumption 3 (Error bound).
-For any $x^{*}\in X^{*}$, $\left\lVert J(x)^{T}F(x)\right\rVert$ provides an error bound in $B(x^{*},\delta)$, i.e., there exists $\omega\in(0,\infty)$ such that
-$\omega\textup{dist}(x,X^{*})\leq\left\lVert J(x)^{T}F(x)\right\rVert,\quad\forall x\in B(x^{*},\delta),$ (10)
+Next, we present a local error bound condition assumed in this work.
+Assumption 2
+For some $x^{*}\in X^{*}$, $\left\lVert J(x)^{T}F(x)\right\rVert$ provides a local error bound at $x^{*}$, i.e., there exists $\delta\in]0,1[$ and $\omega>0$ such that
+$\omega\textup{dist}(x,X^{*})\leq\left\lVert J(x)^{T}F(x)\right\rVert,\quad\forall x\in B(x^{*},\delta),$
 where $\textup{dist}(x,X^{*})=\inf_{z\in X^{*}}{\left\lVert x-z\right\rVert}$.
-Throughout the text, given $x$, we shall denote by $\bar{x}$ an element of $X^{*}$ such that $\|x-\bar{x}\|=\text{dist}(x,X^{*})$. Also, let $x^{*}\in X^{*}$ and $J_{*}=J(x^{*})$.
-From Assumptions 3 and (7), we obtain
-$\omega\text{dist}(x_{k},X^{*})\leq\left\lVert J_{k}^{T}F_{k}\right\rVert\leq L_{3}\text{dist}(x_{k},X^{*}).$ (11)
-The remaining assumption focuses, as outlined in (8), on the terms
+Error bound conditions have been extensively studied and used in convergence analysis of LM methods in the last two decades Yamashita ; fan ; FanYuan ; Fischer2002 ; behling2012 ; behling2013 ; behling2019local ; Fischer2024 .
+For zero residual problems, it is common to use $\|F(x)\|$ as an error bound.
+However, for nonzero residual problems, $F(x)\neq 0,\forall x\in\mathbb{R}^{n}$, and Assumption 2 uses $\|\nabla\phi(x)\|$ as an error bound instead.
+Section 6 presents some examples where such assumption holds.
+See also the discussion and examples in behling2019local .
+It is well-known that Assumption 2 is suitable to deal with nonisolated stationary points and it is weaker than the assumption of the Jacobian having full rank at $x^{*}\in X^{*}$.
+By the way, $J(x^{*})$ having full rank implies Assumption 2.
+From Assumption 2 and (6), we obtain
+$\omega\text{dist}(x,X^{*})\leq\left\lVert J(x)^{T}F(x)\right\rVert\leq L_{3}\text{dist}(x,X^{*}).$ (9)
+The remaining assumptions focus, as outlined in (7), on the terms
 $\left\lVert J(x)^{T}F(z)\right\rVert\text{ and }\left\lVert J(y)^{T}F(z)\right\rVert.$
-These terms play a crucial role in controlling the error of the “incomplete linearization” of the gradient, as expressed in (9[2].
-Assumption 4.
-For every $x\in B(x^{*},\delta)$ and every $z\in X^{*}\cap B(x^{*},\delta)$, the following inequality holds:
-$\|(J(x)-J(z))^{T}F(z)\|\leq C\|x-z\|^{1+r},$ (12)
-with $r\in(0,1]$ and $C\geq 0$.
+These terms play a crucial role in controlling the error, as expressed in (8), of the “incomplete linearization” of the gradient: notice that $J(x)^{T}J(x)$ is used in (8) instead of the Hessian $\nabla^{2}\phi(x)=J(x)^{T}J(x)+S(x)$, with $S(x)=\sum_{i=1}^{m}F_{i}(x)\nabla^{2}F_{i}(x)$, that would appear in a first order Taylor approximation of $\nabla\phi$.
+Assumption 3
+For some $x^{*}\in X^{*}$, for all $x\in B(x^{*},\delta)$ and all $z\in X^{*}\cap B(x^{*},\delta)$, the following inequality holds:
+$\|J(x)^{T}F(z)\|\leq\sigma\|x-z\|,$ (10)
+with $0\leq\sigma<\bar{\sigma}$, where $\bar{\sigma}$ is a positive constant depending on $\omega$, $L_{3}$ and the smallest positive eigenvalue of $J(x^{*})^{T}J(x^{*})$.
+Explicit expressions for $\bar{\sigma}$ will be given ahead.
+Assumption 3 dates back, at least, to the work of Dennis Dennis1977 , and it was also considered in behling2019local where the authors analyze the local convergence of the standard LM (where $L^{T}L=I$) in the case of nonzero residual.
+It is a condition to control the linearization error of the gradient.
+In Dennis1977 , convergence analysis of the classic LMM is given for the nonzero residual case, under the assumption that $J(x^{*})^{T}J(x^{*})$ is nonsingular and $\|J(x)^{T}F(x^{*})\|\leq\sigma\|x-x^{*}\|$, in a neighborhood of $x^{*}\in X^{*}$, for a sufficiently small $\sigma>0$, namely $\sigma<\lambda^{*}:=\lambda_{\min}(J(x^{*})^{T}J(x^{*}))$.
+Here, because we are not requiring $J(x^{*})$ to have full column rank, the set $X^{*}\cap B(x^{*},\delta)$ may be a nonisolated set of stationary points and for this reason Assumption 3 is slightly different from that of Dennis1977 .
+Remark 1
+Observe that
+$\|(J(x)-J(z))^{T}F(z)\|\leq\|J(x)-J(z)\|\|F(z)\|\leq L_{0}\|F(z)\|\|x-z\|.$
+Thus, given a bound $\|F(z)\|\leq\beta$ for every $z\in X^{*}\cap B(x^{*},\delta)$, if $L_{0}\beta<\bar{\sigma}$, then (10) holds with $\sigma=L_{0}\beta$.
+This might occur when the residual is small enough in $X^{*}\cap B(x^{*},\delta)$.
+Assumption 4
+For some $x^{*}\in X^{*}$, for all $x\in B(x^{*},\delta)$ and all $z\in X^{*}\cap B(x^{*},\delta)$, the following inequality holds:
+$\|J(x)^{T}F(z)\|\leq C\|x-z\|^{1+r},$
+with $r\in]0,1]$ and $C\geq 0$.
+Assumption 4 is admittedly stronger than Assumption 3, but it might hold in some specific problems.
+For example, it is clear that such assumption holds when $F$ is linear or $F(z)=0$.
+In behling2019local the authors showed that when Assumption 4 holds, LMM local convergence rate can be superlinear under a suitable choice of the LM parameter.
+In Dennis_Schnabel , it is mentioned that $J(x)^{T}F(z)\approx S(z)(x-z)$ and thus $\sigma$ could be interpreted as a combined measure of nonlinearity and residual size.
+Observe that $S(z)=0$ is also a sufficient (but not necessary) condition to ensure Assumption 4.
+See the examples in Section 6 and in the reference behling2019local .
 3 Auxiliary Results
-In this section, we organize other fundamental results essential for the upcoming analysis.
-We start by mentioning that (2) always has a solution. In fact, (2) corresponds to the normal equation $B^{T}Bd=B^{T}c$ of the augmented system
-$Bd:=\left[\begin{array}[]{c}J_{k}\\ \sqrt{\lambda_{k}}L\end{array}\right]d=-\left[\begin{array}[]{c}F_{k}\\ 0\end{array}\right]=:c,$
-and the normal equation always has a solution because $B^{T}c$ lies in the range of $B^{T}B$.
-But it is Assumption 1 that ensures the solution of (2) is unique.
-Indeed, it is easy to see that condition (4) at $x_{k}$ implies that $J_{k}^{T}J_{k}+\lambda_{k}L^{T}L$ is positive definite (and thus nonsingular).
-Next result will be important when the rank of $J(x)^{T}J(x)$ is constant in a neighborhood of $x^{*}$.
-Lemma 2.
-[2, Corollary 3.1] Given $\kappa>1$, if
-$\text{rank}(J(x)^{T}J(x))=\text{rank}(J_{*}^{T}J_{*})=q\geq 1,$
-and
+In this section, we organize other fundamental results for the upcoming analysis.
+We start with a perturbation lemma that will be important when the rank of $J(x)^{T}J(x)$ is constant in a neighborhood of $x^{*}$.
+Lemma 2
+(behling2019local, , Corollary 3.1) Given $\kappa>1$, if $\text{rank}(J(x)^{T}J(x))=\text{rank}(J_{*}^{T}J_{*})=q\geq 1$, and
 $\left\lVert J(x)^{T}J(x)-J_{*}^{T}J_{*}\right\rVert\leq\Big{(}1-\frac{1}{\kappa}\Big{)}\frac{1}{\left\lVert(J_{*}^{T}J_{*})^{+}\right\rVert},$ (11)
@@
 $\left\lVert(J(x)^{T}J(x))^{+}\right\rVert\leq\kappa\left\lVert(J_{*}^{T}J_{*})^{+}\right\rVert.$
-Now, we will show some results on the Generalized Singular Value Decomposition (GSVD), an important tool in theoretical analysis of the LMMSS direction.
-Theorem 1.
-(GSVD) [9, p.
-22] Consider the pair $(A,L)$, where $A\in\mathbb{R}^{m\times n}$, $L\in\mathbb{R}^{p\times n}$, $m\geq n\geq p$, $rank(L)=p$, and $\mathcal{N}(A)\cap\mathcal{N}(L)=\{0\}$.
+Now, we will recall some results on the Generalized Singular Value Decomposition (GSVD), an important tool in theoretical analysis of the LMMSS direction.
+Theorem 1
+(GSVD) (Hansen1998, , p.
+22) Consider the pair $(A,L)$, where $A\in\mathbb{R}^{m\times n}$, $L\in\mathbb{R}^{p\times n}$, $m\geq n\geq p$, $rank(L)=p$, and $\mathcal{N}(A)\cap\mathcal{N}(L)=\{0\}$.
 Then, there exist matrices $U\in\mathbb{R}^{m\times n}$ and $V\in\mathbb{R}^{p\times p}$ with orthonormal columns and a nonsingular matrix $X\in\mathbb{R}^{n\times n}$ such that
-$A=U\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right]X^{-1}\quad\text{and}\quad L=V\left[\begin{array}[]{cc}M&0\end{array}\right]X^{-1},$ (14)
+$A=U\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right]X^{-1}\quad\text{and}\quad L=V\left[\begin{array}[]{cc}M&0\end{array}\right]X^{-1},$
 with $\Sigma$ and $M$ being the following diagonal matrices:
-$\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{p})\in\mathbb{R}^{p\times p}\quad\text{and}\quad M=\text{diag}(\mu_{1},\dots,\mu_{p})\in\mathbb{R}^{p\times p}.$ (15)
-Moreover, the elements of $\Sigma$ and $M$ are non-negative, ordered as follows:
-$0\leq\sigma_{1}\leq\dots\leq\sigma_{p}\leq 1\quad\text{and}\quad 1\geq\mu_{1}\geq\dots\geq\mu_{p}>0,$ (16)
+$\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{p})\in\mathbb{R}^{p\times p}\quad\text{and}\quad M=\text{diag}(\mu_{1},\dots,\mu_{p})\in\mathbb{R}^{p\times p}.$
+Moreover, the elements of $\Sigma$ and $M$ are nonnegative, ordered as follows:
+$0\leq\sigma_{1}\leq\dots\leq\sigma_{p}\leq 1\quad\text{and}\quad 1\geq\mu_{1}\geq\dots\geq\mu_{p}>0,$
 and normalized by the relation $\sigma_{i}^{2}+\mu_{i}^{2}=1$, for $i=1,\dots,p$. We call the generalized singular value of the pair $(A,L)$ the ratio
 $\gamma_{i}=\frac{\sigma_{i}}{\mu_{i}},\quad i=1,\dots,p.$
-Remark 1.
+Remark 2
 Since $X^{-1}$ is nonsingular and $U$ has orthonormal columns we have
@@
 with $\Gamma_{k}:=(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}.$
-Next result from [5] gives an upper bound on $\|X_{k}\|$ depending on $\gamma$ from Assumption 1.
-Lemma 3.
-[5, Lemma 2.2] Consider $\{x_{k}\}$ the sequence generated by LMMSS along with the GSVD of the pair $(J_{k},L)$. Then, $\forall k\in\mathbb{N}$, $\|X_{k}\|\leq 1/\sqrt{\gamma}$.
-Remark 2.
+The following result provides bounds for the matrix $X_{k}$.
+Lemma 3
+(Boos2024, , Lemma 2.2) Consider $\{x_{k}\}$ the sequence generated by LMMSS along with the GSVD of the pair $(J_{k},L)$.
+If $\{x_{k}\}\subset\Omega$ where the uniform completeness condition holds, then for each $k\in\mathbb{N}$, we have
+$\|X_{k}\|\leq\dfrac{1}{\sqrt{\gamma}}.$ (15)
+Remark 3
 By the GSVD for the pair $(J_{k},L)$, it follows that
@@
 Therefore, we can bound $\|\Gamma_{k}\|$ in terms of the generalized singular values $\gamma_{i,k}$ and the parameter $\lambda_{k}$. To this end, we can use the following lemma.
-Lemma 4.
+Lemma 4
 For the function
@@
   • (a)
-    For $\lambda\in(0,1)$, the function $\psi(\gamma,\lambda)$ has a unique maximum value attained at
+    For $\lambda\in]0,1[$, the function $\psi(\gamma,\lambda)$ has a unique maximum value attained at
     $\gamma_{\max}=0\quad\text{ where }\quad\psi(\gamma_{\max},\lambda)=\max_{\gamma\geq 0}\psi(\gamma,\lambda)=\frac{1}{\lambda}.$
   • (b)
-    For a fixed $\lambda\in[1,+\infty)$, the function $\psi(\gamma,\lambda)$ is non-decreasing and upper-bounded. More precisely, $\psi(\gamma,\lambda)\leq 1,\quad\forall\gamma\geq 0.$
-Proof.
+    For a fixed $\lambda\in[1,+\infty)$, the function $\psi(\gamma,\lambda)$ is nondecreasing and upper-bounded. More precisely,
+    $\psi(\gamma,\lambda)\leq 1,\quad\forall\gamma\geq 0.$
+Proof
 First, note that for any fixed $\lambda$, we have
-$\lim_{\gamma\rightarrow\infty}\psi(\gamma,\lambda)=1.$ (21)
+$\lim_{\gamma\rightarrow+\infty}\psi(\gamma,\lambda)=1=\lim_{\gamma\rightarrow-\infty}\psi(\gamma,\lambda).$ (17)
 Now, given $\lambda>0$, observe that
@@
   • (a)
-    Consider $\lambda\in(0,1)$ fixed.
-As the maximizers of $\psi(\gamma,\lambda)$, for $\gamma>0$, must satisfy $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)=0,$ from (22), we have $02(\lambda-1)\gamma=$ 0, thus $\gamma_{\max}=0$.
+    Consider $\lambda\in]0,1[$ fixed. As the maximizers of $\psi(\gamma,\lambda)$, for $\gamma\in\mathbb{R}$, must satisfy $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)=0,$ from (18), we have
+    $2(\lambda-1)\gamma=0\quad\Rightarrow\quad\gamma_{\max}=0,$
     Substituting $\gamma_{\max}$ into $\psi(\gamma,\lambda)$, we obtain
-    $\psi(\gamma_{\max},\lambda)=1/2.$
+    $\psi(\gamma_{\max},\lambda)=\frac{1}{\lambda}.$
   • (b)
-    Now, for fixed $\lambda\in[1,+\infty)$, from (22), we have $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)>0$ for $\gamma\geq 0$.
-Therefore, $\psi$ is increasing for $\gamma\in[0,+\infty)$, with $\psi(0,\lambda)=0$.
+    Now, for fixed $\lambda\in[1,+\infty)$, from (18), we have $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)>0$ for $\gamma>0$.
+Therefore, $\psi$ is increasing for $\gamma\in[0,+\infty)$, with $\psi(0,\lambda)=1/\lambda\leq 1$.
 Then, from (17), we conclude that $\psi(\gamma,\lambda)\leq 1$, for all $\gamma\geq 0$.
+∎
 We are now ready to prove two key lemmas for the local convergence analysis.
-They show that the norm of LMMSS direction $d_{k}$ is ${\cal O}(\textup{dist}(x_{k},X^{*}))$, considering two distinct cases.
-Lemma 5 addresses the scenario where the rank of the Jacobian around the solution (stationary-point) is constant, while Lemma 6 deals with the case of diminishing rank.
-Lemma 5.
-Suppose that Assumptions 1-3 are valid in $B(x^{*},\delta)$, and
+They show that the norm of LMMSS direction $d_{k}$ is $O(\textup{dist}(x_{k},X^{*}))$, considering two distinct cases.
+Lemma 5 addresses the scenario where the rank of the Jacobian around a stationary-point is constant, while Lemma 6 deals with the case when the rank decreases [^2] Observe that the rank function is lower semicontinuous thus, when a sequence $x_{k}$ in a neighborhood of $x^{*}\in X^{*}$ approaches $x^{*}$, the rank of the Jacobian $J(x_{k})$ either remains constant or decreases, these are the only two possibilities..
+Lemma 5
+Suppose that Assumption 1 is valid in $B(x^{*},\delta)$, that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$, and
 $rank(J(x)^{T}J(x))=rank(J(x^{*})^{T}J(x^{*}))=q\geq 1$
 for all $x\in B(x^{*},\delta)$. If $x_{k}\in B(x^{*},\delta)$, $\lambda_{k}>0$ and $J_{k}^{T}F_{k}\neq 0$, then there exists $c_{1}>0$ such that
-$\|d_{k}\|\leq c_{1}\textup{dist}(x_{k},X^{*}).$ (23)
-Proof.
+$\|d_{k}\|\leq c_{1}\textup{dist}(x_{k},X^{*}).$
+Proof
 Recall that $J_{k}^{T}F_{k}\in\mathcal{R}(J_{k}^{T})=\mathcal{N}(J_{k})^{\perp}=\mathcal{N}(J_{k}^{T}J_{k})^{\perp}$, and
@@
 which implies that $\left\lVert J_{k}d_{k}\right\rVert^{2}+\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=-(J_{k}d_{k})^{T}F_{k}$, thus, if we assume that $d_{k}\in\mathcal{N}(J_{k})$, we conclude that $\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=0$.
-But $\left\lVert Ld_{k}\right\rVert>0$, given that, by Assumption 1, $\mathcal{N}(J_{k})\cap\mathcal{N}(L)=0$, and also $\lambda_{k}\neq 0$, and we have a contradiction.
+But $\left\lVert Ld_{k}\right\rVert>0$, given that, by condition 4, $\mathcal{N}(J_{k})\cap\mathcal{N}(L)=0$, and also $\lambda_{k}\neq 0$, and we have a contradiction.
 We can write $d_{k}=d_{N}+d_{R},$ with $d_{N}\in\mathcal{N}(J_{k})$ and $d_{R}\in\mathcal{N}(J_{k})^{\perp}=\mathcal{R}(J_{k}^{T})$. Since $d_{k}\notin\mathcal{N}(J_{k})$, we have $d_{R}\neq 0$.
 Thus, from (19)
-$-J_{k}^{T}F_{k}$ $=(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}=(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)(d_{N}+d_{R})$ (25)
+$-J_{k}^{T}F_{k}$ $=(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}$ (20)
+$=(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)(d_{N}+d_{R})$
 $=J_{k}^{T}J_{k}d_{N}+J_{k}^{T}J_{k}d_{R}+\lambda_{k}L^{T}Ld_{N}+\lambda_{k}L^{T}Ld_{R}$
@@
 Let $J_{k}=U\Sigma V^{T}$ be the SVD for $J_{k}$.
-As $d_{R}\in\mathcal{N}(J_{k})^{\perp}$, $d_{R}$ can be written as a linear combination of the columns of $V$ corresponding to the non-zero singular values of $J_{k}$.
+As $d_{R}\in\mathcal{N}(J_{k})^{\perp}$, $d_{R}$ can be written as a linear combination of the columns of $V$ corresponding to the nonzero singular values of $J_{k}$.
 Therefore,
@@
 Now let us show that $d_{R}\notin\mathcal{N}(L)$.
-Suppose $d_{R}\in\mathcal{N}(L)$, from (26), we get $\left\lVert Ld_{N}\right\rVert^{2}=-d_{N}^{T}L^{T}Ld_{R}=0,$ implying that $d_{N}\in\mathcal{N}(L)$ and $d_{N}\in\mathcal{N}(J_{k})$, which is a contradiction with Assumption 1.
-By Assumption 1, there exists $\gamma$ such that $\gamma\left\lVert d_{N}\right\rVert^{2}\leq\left\lVert J_{k}d_{N}\right\rVert^{2}+\left\lVert Ld_{N}\right\rVert^{2}$, which implies
+Suppose $d_{R}\in\mathcal{N}(L)$, from (21), we get $\left\lVert Ld_{N}\right\rVert^{2}=-d_{N}^{T}L^{T}Ld_{R}=0,$ implying that $d_{N}\in\mathcal{N}(L)$ and $d_{N}\in\mathcal{N}(J_{k})$, which is a contradiction with condition (5).
+By condition (4), there exists $\gamma>0$ such that
+$\gamma\left\lVert d_{N}\right\rVert^{2}\leq\left\lVert J_{k}d_{N}\right\rVert^{2}+\left\lVert Ld_{N}\right\rVert^{2},$
+which implies
 $\left\lVert d_{N}\right\rVert^{2}\leq\frac{1}{\gamma}\left\lVert Ld_{N}\right\rVert^{2}.$ (26)
 Therefore, from (22), (25), (26), and Assumption 1,
-$\left\lVert d_{k}\right\rVert^{2}$ $=\left\lVert d_{N}+d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+2d_{N}^{T}d_{R}+\left\lVert d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}$ (32)
+$\left\lVert d_{k}\right\rVert^{2}$ $=\left\lVert d_{N}+d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+2d_{N}^{T}d_{R}+\left\lVert d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}$
 $\leq\frac{1}{\gamma}\left\lVert Ld_{N}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}\leq\frac{1}{\gamma}\left\lVert Ld_{R}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}$
-$\leq\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}\right\rVert^{2}=\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}-J_{*}^{T}F_{*}\right\rVert^{2}$
+$\leq\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}\right\rVert^{2}$
+$=\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}-J_{*}^{T}F_{*}\right\rVert^{2}$
 $\leq\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}L_{3}^{2}\textup{dist}^{2}(x_{k},X^{*}).$
-Since $J(x)$ is continuous, it is clear that for a sufficiently small $\delta>0$, condition (13) is satisfied for all $x_{k}\in B(x^{*},\delta)$. Finally, we have
+Since $J(x_{k})^{T}J(x_{k})$ is continuous, it is clear that for a sufficiently small $\delta>0$, condition (11) is satisfied for all $x_{k}\in B(x^{*},\delta)$ and applying Lemma 2 we have
 $\left\lVert d_{k}\right\rVert\leq\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{q}^{*}}L_{3}\textup{dist}(x_{k},X^{*}),$
@@
 ∎
-Lemma 6.
-Suppose that Assumptions 13 are valid in $B(x^{*},\delta)$, for some $\delta>0$.
-If $x_{k}\in B(x^{*},\delta)$, $rank(J_{k})=\ell\geq rank(J_{*})=q\geq 1$, and Assumption 4 is satisfied, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r},$ with $r\in(0,1]$, then there exists $c_{1}>0$ such that $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$.
-Proof.
-Consider the GSVD for the pair $(J_{k},L)$. In this case where $rank(J_{k})=\ell,rank(J_{*})=q\geq 1$ and $rank(L)=p$ with $n-q\leq p$ and $n-\ell\leq p$ (see Remark 1), we remark that:
+Lemma 6
+Suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$, for some $\delta\in]0,1[$ and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$.
+If $rank(J_{k})=\ell\geq rank(J_{*})=q\geq 1$, and
+(a) $x_{k}\in B(x^{*},\delta)\setminus X^{*}$, Assumption 4 is satisfied with $r\in]0,1]$, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, or,
+(b) $x_{k}\in B(x^{*},\bar{\delta})\setminus X^{*}$, Assumption 3 is satisfied with $\sigma<(\sigma_{\min}^{*})^{2}$, where $\sigma_{\min}^{*}$ is the smallest positive singular value of $J(x^{*})$ and
+$L_{4}\|x_{k}-\bar{x}_{k}\|+\sigma\leq\lambda_{k}\leq\theta\left(L_{4}\|x_{k}-\bar{x}_{k}\|+\sigma\right),$ (27)
+where $\theta>1$ is such that $(\sigma_{\min}^{*})^{2}>\theta\sigma$ and $\bar{\delta}=\min\left\{\delta,\frac{(\sigma_{\min}^{*})^{2}-\theta\sigma}{\theta L_{4}}\right\}$,
+then there exists $c_{1}>0$ such that $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$.
+Proof
+Consider the GSVD for the pair $(J_{k},L)$. In this case where $rank(J_{k})=\ell\geq q=rank(J_{*})\geq 1$ and $rank(L)=p$ with $n-q\leq p$ and $n-\ell\leq p$ (see Remark 2), we observe that:
   • For $1\leq i<n-\ell$, we have $\sigma_{i,k}$=0 and $\mu_{i,k}=1.$ This occurs because of the fact that $rank(J_{k})=\ell$ and the relationship between the singular values of the GSVD (see Theorem 1).
@@
   • For $n-\ell\leq i<n-q$, if $\sigma_{i,k}\rightarrow 0$, then $\mu_{i,k}\rightarrow 1$, and we have $(\Gamma_{k})_{ii}\rightarrow\frac{1}{\lambda_{k}}$.
-  • For $n-q\leq i\leq p$, then $(\Gamma_{k})_{ii}=\frac{1}{\sigma_{i,k}+\lambda_{k}\mu_{i,k}}<\frac{1}{\sigma_{i,k}}\leq\frac{1}{\sigma^{*}_{\text{min}}},$ where $\sigma^{*}_{\text{min}}$ is the smallest positive singular value of $J(x^{*}).$
-Note that, for $x_{k}\in B(x^{*},\delta)$ and due to the continuity of $\mu_{i,k}$, in the case where $n-\ell\leq i<n-q$ we can assume that $\mu_{i,k}>\frac{1}{2}$ and thus, for $\delta$ sufficiently small, we have $(\Gamma_{k})_{ii}\leq\frac{1}{4\lambda_{k}}$.
-Therefore,
-$\left\lVert\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]\right\rVert\leq\max{\Big{[}\frac{1}{\lambda_{k}},\frac{1}{4\lambda_{k}},\frac{1}{\sigma^{*}_{\text{min}}},1\Big{]}}\leq\frac{1}{\lambda_{k}},$ (33)
-as $x_{k}\in B(x^{*},\delta)$ and $\left\lVert J_{k}^{T}F_{k}\right\rVert<1.$
-From (19) and (33), we obtain
+  • For $n-q\leq i\leq p$, then $(\Gamma_{k})_{ii}=\frac{1}{\sigma_{i,k}^{2}+\lambda_{k}\mu_{i,k}^{2}}<\frac{1}{\sigma_{i,k}^{2}}\leq\frac{\kappa}{(\sigma^{*}_{\text{min}})^{2}},$ where $\sigma^{*}_{\text{min}}$ is the smallest positive singular value of $J(x^{*})$ and $\kappa>1$ (cf.
+Lemma 2).
+Note that, for $x_{k}\in B(x^{*},\delta)$ and due to the continuity of $\mu_{i,k}$, in the case where $n-\ell\leq i<n-q$ we can assume that $\mu_{i,k}>\frac{1}{\sqrt{\kappa}}$ and thus, for $\delta$ sufficiently small, we have $(\Gamma_{k})_{ii}\leq\frac{\kappa}{\lambda_{k}}$.
+(a) If $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, for $\delta$ small enough, $\lambda_{k}<\min\{(\sigma_{\min}^{*})^{2}/\kappa,1\}$, and we obtain
+$\left\lVert\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]\right\rVert\leq\max{\Big{[}\frac{1}{\lambda_{k}},\frac{\kappa}{\lambda_{k}},\frac{\kappa}{(\sigma^{*}_{\text{min}})^{2}},1\Big{]}}\leq\frac{\kappa}{\lambda_{k}}.$ (28)
+From (14) and (28), we get
 $\left\lVert d_{k}\right\rVert$ $=\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}{J_{k}}^{T}F_{k}\right\rVert$ (29)
-$=\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}\Big{(}{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})\Big{)}\right\rVert$
-$\leq\left\lVert X_{k}\right\rVert^{2}\frac{1}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$
+$=\left\|X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}\Big{(}{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-\right.$
+$\qquad\left.-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})+J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\Big{)}\right\|$
+$\leq\left\lVert X_{k}\right\rVert^{2}\frac{\kappa}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$
 $+\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert.$
@@
 Therefore, using (14), (30), (8), (16) and Lemma 3 in (29), we have
-$\left\lVert d_{k}\right\rVert$ $\leq\left\lVert X_{k}\right\rVert^{2}\frac{1}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$ (36)
+$\left\lVert d_{k}\right\rVert$ $\leq\left\lVert X_{k}\right\rVert^{2}\frac{\kappa}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$ (31)
 $+\left\lVert X_{k}\right\rVert\left\lVert X_{k}^{-1}\right\rVert\left\lVert x_{k}-\bar{x}_{k}\right\rVert$
-$\leq\frac{1}{\gamma\lambda_{k}}L_{4}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{2}+\frac{1}{\gamma\lambda_{k}}\left\lVert J_{k}^{T}F(\bar{x}_{k})\right\rVert+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
-Finally, from Assumption 4 and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, the result follows. ∎
+$\leq\frac{\kappa}{\gamma\lambda_{k}}L_{4}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{2}+\frac{\kappa}{\gamma\lambda_{k}}\left\lVert J_{k}^{T}F(\bar{x}_{k})\right\rVert+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
+From Assumptions 2 and 4, the choice $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ and (9), we obtain from (31) that
+$\left\lVert d_{k}\right\rVert$ $\leq\left(\frac{\kappa L_{4}}{\gamma\omega^{r}}+\frac{\kappa C}{\gamma\omega^{r}}+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\right)\left\lVert x_{k}-\bar{x}_{k}\right\rVert,$
+which concludes the proof for (a) with $c_{1}=\kappa(L_{4}+C)/\gamma\omega^{r}+\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}/\sqrt{\gamma}$.
+(b) Let us denote
+$\Theta:=\max{\Big{[}\frac{1}{\lambda_{k}},\frac{\kappa}{\lambda_{k}},\frac{\kappa}{(\sigma^{*}_{\text{min}})^{2}},1\Big{]}},$
+Instead of (31), we may write
+$\|d_{k}\|\leq\frac{\Theta}{\gamma}L_{4}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{2}+\frac{\Theta}{\gamma}\left\lVert J_{k}^{T}F(\bar{x}_{k})\right\rVert+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
+Then, from Assumption 3, we have
+$\left\lVert d_{k}\right\rVert$ $\leq\left(\frac{\Theta}{\gamma}(L_{4}\|x_{k}-\bar{x}_{k}\|+\sigma)+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\right)\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
+Since, $\|x_{k}-\bar{x}_{k}\|\leq\delta\leq\frac{(\sigma_{\min}^{*})^{2}-\theta\sigma}{\theta L_{4}}$, it follows that $\lambda_{k}\leq(\sigma_{\min}^{*})^{2}$.
+Then, $\kappa/\lambda_{k}\geq\kappa/(\sigma_{\min}^{*})^{2}$.
+Clearly, $\kappa/\lambda_{k}>1/\lambda_{k}$ because $\kappa>1$.
+Hence, $\Theta=\max\{1,\kappa/\lambda_{k}\}$.
+Now, assume $\lambda_{k}$ from (27). In case $\Theta=\kappa/\lambda_{k}$, the above inequality becomes
+$\left\lVert d_{k}\right\rVert\leq\left(\frac{\kappa}{\gamma}+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\right)\left\lVert x_{k}-\bar{x}_{k}\right\rVert,$
+If $\Theta=1$, from the upper bound on $\delta$ in (b) we obtain
+$\left\lVert d_{k}\right\rVert\leq\left(\frac{(\sigma_{\min}^{*})^{2}}{\theta\gamma}+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\right)\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
+Therefore, we get the proof for (b) with $c_{1}=(1/\gamma)\max\{\kappa,(\sigma_{\min}^{*})^{2}/\theta\}+\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}/\sqrt{\gamma}$. ∎
 4 Local convergence
-Lemmas 5 and 6 showed that $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, under specific assumptions.
+Now we focus on the local convergence of the “pure” LMMSS iteration, i.e iteration (2)–(3) with $\alpha_{k}=1$.
+We assume the initial point $x_{0}$ is in a neighborhood of a (possibly nonisolated) stationary point $x^{*}\in X^{*}$.
+Lemmas 5 and 6 showed that $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$.
 As we shall see in this section, such inequality is key for the local convergence analysis under an error bound condition (Assumption 2).
+As we mentioned in the introduction, the analysis developed in this section is based on the work behling2019local and extends it to handle singular scaling in LM methods.
 Apart from the next lemma which is an intermediate result, the remaining of this section is organized in subsections according to whether the Jacobian rank near the solution set is constant or not.
-Throughout this section, we consider the LMMSS iteration (2)–(3) with $\alpha_{k}=1$ for every $k$, that is, the “pure” LMMSS.
-Lemma 7.
-Let $\{x_{k}\}$ be the sequence generated by the LMMSS (with $\alpha_{k}=1$) and suppose Assumptions 1-3 hold.
+Lemma 7
+Let $x^{*}\in X^{*}$ and $\{x_{k}\}$ be the sequence generated by the LMMSS (with $\alpha_{k}=1$).
+Suppose Assumptions 1 and 2 holds in $B(x^{*},\delta)$, for some $\delta\in]0,1[$ and that the uniform completeness condition is valid in $B(x^{*},\delta)\setminus X^{*}$.
 If $x_{k+1},x_{k}\in B(x^{*},\delta)$ and $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, then
@@
 where $L_{5}:=L_{0}L_{2}(1+c_{1})(2+c_{1})$.
-Proof.
+Proof
 For any $x,y\in\mathbb{R}^{n}$, the reverse triangle inequality implies
@@
 Thus, using (35) and (36) in (34) yields
-$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq L_{4}\|d_{k}\|^{2}+L_{0}L_{2}\|x_{k}-\bar{x}_{k}\|\|x_{k+1}-\bar{x}_{k}\|+\|J_{k}^{T}F(\bar{x}_{k})\|$ (42)
+$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq L_{4}\|d_{k}\|^{2}+L_{0}L_{2}\|x_{k}-\bar{x}_{k}\|\|x_{k+1}-\bar{x}_{k}\|+\|J_{k}^{T}F(\bar{x}_{k})\|$
 $+L_{0}L_{2}\|x_{k+1}-\bar{x}_{k}\|^{2}+\|J_{k+1}F(\bar{x}_{k})\|+\lambda_{k}\left\lVert L\right\rVert^{2}\left\lVert d_{k}\right\rVert$
@@
 $+L_{0}L_{2}(1+c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+L_{0}L_{2}(1+c_{1})^{2}\|x_{k}-\bar{x}_{k}\|^{2}$
-$+\|J_{k}^{T}F(\bar{x}_{k})\|+\|J_{k+1}^{T}F(\bar{x}_{k})\|.$
+$+\|J_{k}^{T}F(\bar{x}_{k})\|+\|J_{k+1}^{T}F(\bar{x}_{k})\|,$
+concluding the proof. ∎
 4.1 Constant Rank
 In this section, we assume that $1\leq\text{rank}(J(x^{*}))=q\leq\min\{n,m\}$, and $\text{rank}(J(x))=\text{rank}(J(x^{*}))$ for each $x\in B(x^{*},\delta)$.
-From Lemma 5, we recall that $\left\|d_{k}\right\|\leq c_{1}\text{dist}(x_{k},X^{*})$, with $c_{1}=\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{p}^{*}}L_{3}$, provided that $\lambda_{k}>0,$ for $x_{k}\neq x^{*}$.
+From Lemma 5, we recall that $\left\|d_{k}\right\|\leq c_{1}\text{dist}(x_{k},X^{*})$, with $c_{1}=\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{p}^{*}}L_{3}$.
 Next, we present two lemmas that will aid in proving the convergence theorem under Assumption 3.
-Lemma 8.
-Suppose that Assumptions 1-3 are valid, and $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$.
+Lemma 8
+Let $x^{*}\in X^{*}$, suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$, and the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$.
+Assume $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$, that Assumption 3 is verified with $\sigma<\bar{\sigma}:=\omega/(2+c_{1})$, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|$.
+If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
+$\varepsilon\leq\frac{\eta\omega-(2+c_{1})\sigma}{L_{6}},$
+with $\eta\in\left]\frac{\sigma(2+c_{1})}{\omega},1\right[$ and $L_{6}=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$, then
+$\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
+Proof
+Since $J(x)$ has constant rank in $B(x^{*},\delta)$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, and Assumption 3 is satisfied, from (32) and (6), we have
+$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||$
+$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||$
+$\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{2}$
+$+\sigma\|x_{k}-\bar{x}_{k}\|+\sigma\|x_{k+1}-\bar{x}_{k}\|$
+$\leq(L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+\sigma\|x_{k}-\bar{x}_{k}\|$
+$+(1+c_{1})\sigma\|x_{k}-\bar{x}_{k}\|$
+$\leq(L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+(2+c_{1})\sigma\|x_{k}-\bar{x}_{k}\|$
+$\leq\Big{[}(L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1})\varepsilon+(2+c_{1})\sigma\Big{]}\|x_{k}-\bar{x}_{k}\|.$
+As $\sigma<\frac{\omega}{2+c_{1}}$, for any $\eta\in(\frac{\sigma(2+c_{1})}{\omega},1)$, we have $\eta\omega-(2+c_{1})\sigma>0$.
+Denoting $L_{6}:=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$, for $\varepsilon\leq\frac{\eta\omega-(2+c_{1})\sigma}{L_{6}}$, we obtain $\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*})$.
+∎
+Lemma 9
+(behling2019local, , Lemma 4.3) Suppose that the assumptions of Lemma 8 are satisfied, and consider
+$\varepsilon=\min\left\{\frac{\frac{\delta}{2}}{1+\frac{c_{1}}{1-\eta}},\frac{\eta\omega-(2+c_{1})\sigma}{L_{6}}\right\}.$ (37)
+If $x_{0}\in B(x^{*},\varepsilon)$, then $x_{k+1}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{k},X^{*})\leq\varepsilon$, for every $k\in\mathbb{N}$.
+Proof
+The proof is done by induction. For $k=0$, we have $\textup{dist}(x_{0},X^{*})\leq\|x_{0}-x^{*}\|\leq\varepsilon$, thus
+$\|x_{1}-x^{*}\|\leq\|x_{1}-x_{0}\|+\|x_{0}-x^{*}\|\leq\|d_{0}\|+\varepsilon\leq c_{1}\textup{dist}(x_{0},X^{*})+\varepsilon\leq(1+c_{1})\varepsilon<\frac{\delta}{2},$
+and by Lemma  8, $\textup{dist}(x_{1},X^{*})\leq\eta\textup{dist}(x_{0},X^{*})\leq\varepsilon$.
+From the induction hypothesis for $i\leq k$, $x_{i}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{i-1},X^{*})\leq\varepsilon$, for $i=1,...,k$.
+Then, by successively applying Lemma 8 to $x_{i-1}$ and $x_{i}$, for $i=1,...,k$, we have
+$\textup{dist}(x_{k},X^{*})\leq\eta\textup{dist}(x_{k-1},X^{*})\leq\cdots\leq\eta^{k}\textup{dist}(x_{0},X^{*})\leq\eta^{k}\varepsilon<\varepsilon.$
+Finally, using Lemma 5,
+$\|x_{k+1}-x^{*}\|$ $\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}\|d_{i}\|\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}c_{1}\textup{dist}(x_{i},X^{*})$
+$\leq(1+c_{1})\varepsilon+c_{1}\varepsilon\sum_{i=1}^{\infty}\eta^{i}=\Big{(}1+\frac{c_{1}}{1-\eta}\Big{)}\varepsilon\leq\frac{\delta}{2},$
+which completes the proof. ∎
+Theorem 2
+Let $x^{*}\in X^{*}$, suppose that Assumptions 1, 2 and 3 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$ and $\bar{\sigma}=\omega/(2+c_{1})$, $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$.
+Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ for all $k$, and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (37).
+Then, $\{\textup{dist}(x_{k},X^{*})\}$ converges linearly to zero.
+Moreover, the sequence $\{x_{k}\}$ converges to some $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
+Proof
+The linearly convergence of $\textup{dist}(x_{k},X^{*})$ to zero follows directly from Lemma 8 and Lemma 9.
+Once $\textup{dist}(x_{k},X^{*})$ converges to zero and $x_{k}\in B(x^{*},\frac{\delta}{2})$ for every $k$, it remains to show that $\{x_{k}\}$ converges. From Lemma 5 and Lemma 8, we have
+$\|d_{k}\|\leq c_{1}\text{dist}(x_{k},X^{*})\leq c_{1}\eta^{k}\text{dist}(x_{0},X^{*})\leq c_{1}\varepsilon\eta^{k},$
+for every $k\geq 1$. Thus, for any positive integers $\ell$ and $q$, with $\ell\geq q$,
+$\|x_{\ell}-x_{q}\|\leq\sum_{i=q}^{\ell-1}\|d_{i}\|\leq\sum_{i=q}^{\infty}\|d_{i}\|\leq c_{1}\varepsilon\sum_{i=q}^{\infty}\eta^{i}\leq\frac{c_{1}\varepsilon}{1-\eta},$
+implying that $\{x_{k}\}\subset\mathbb{R}^{n}$ is a Cauchy sequence and, therefore, converges. ∎
+Now we demonstrate that under Assumption 4, which is stronger than Assumption 3, the local convergence can be superlinear.
+In the lemmata and theorems that follows we shall suppress algebraic manipulations that are similar to the results above.
+Lemma 10
+Let $x^{*}\in X^{*}$, suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$, $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$.
 Additionally, assume that Assumption 4 is verified, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|$.
 If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, with
-$\varepsilon\leq\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r},$ (43)
-with $\eta\in(0,1)$ and $L_{6}=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$, then
+$\varepsilon\leq\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r},$
+with $\eta\in]0,1[$, $\tilde{C}=L_{6}+(1+(1+c_{1})^{1+r})C$ and $L_{6}$ as in Lemma 8, then
 $\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
-Proof.
-We know that $J(x)$ has constant rank in $B(x^{*},\delta)$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, and Assumptions 4 is satisfied. Thus, from (37) and (7), we have
-$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||$ (44)
-$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||$
-$\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{2}$
+Proof
+Similarly to the proof of Lemma 8, we have
+$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{2}$ (38)
 $+C\|x_{k}-\bar{x}_{k}\|^{1+r}+C\|x_{k+1}-\bar{x}_{k}\|^{1+r}$
@@
 $+C(1+c_{1})^{1+r}\|x_{k}-\bar{x}_{k}\|^{1+r}$
-$=L_{6}\|x_{k}-\bar{x}_{k}\|^{2}+(1+(1+c_{1})^{1+r})C\|x_{k}-\bar{x}_{k}\|^{1+r}$
+$=:L_{6}\|x_{k}-\bar{x}_{k}\|^{2}+(1+(1+c_{1})^{1+r})C\|x_{k}-\bar{x}_{k}\|^{1+r}$
 $\leq L_{6}\|x_{k}-\bar{x}_{k}\|^{1+r}+(1+(1+c_{1})^{1+r})C\|x_{k}-\bar{x}_{k}\|^{1+r}$
@@
 and, given $\eta\in]0,1[$, for $\varepsilon\leq\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r}$ we obtain
-$\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*}).$
-Lemma 9.
+$\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*}),$
+which concludes the proof. ∎
+Lemma 11
 Suppose that the assumptions of Lemma 10 are satisfied, and
@@
 If $x_{0}\in B(x^{*},\varepsilon)$, then $x_{k+1}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{k},X^{*})\leq\varepsilon$, for every $k\in\mathbb{N}$.
-Proof.
-The proof is done by induction. For $k=0$, we have $\textup{dist}(x_{0},X^{*})\leq\|x_{0}-x^{*}\|\leq\varepsilon$ and
-$\|x_{1}-x^{*}\|\leq\|x_{1}-x_{0}\|+\|x_{0}-x^{*}\|\leq\|d_{0}\|+\varepsilon\leq c_{1}\textup{dist}(x_{0},X^{*})+\varepsilon\leq(1+c_{1})\varepsilon<\frac{\delta}{2}.$
-For $k\geq 1$, assume that $x_{i}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{i-1},X^{*})\leq\varepsilon$ for $i=1,...,k$. Then, from Lemma 8 applied to $x_{k-1}$ and $x_{k}$, we obtain
-$\|x_{k}-\bar{x}_{k}\|=\textup{dist}(x_{k},X^{*})\leq\eta\textup{dist}(x_{k-1},X^{*})\leq\eta\varepsilon<\varepsilon.$
-Moreover, by Lemma 8 and the induction hypothesis for $i\leq k$, we have
-$\textup{dist}(x_{i},X^{*})\leq\eta\textup{dist}(x_{i-1},X^{*})\leq\cdots\leq\eta^{i}\textup{dist}(x_{0},X^{*})\leq\eta^{i}\varepsilon<\varepsilon.$
-Thus, using Lemma 5,
-$\|x_{k+1}-x^{*}\|$ $\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}\|d_{i}\|\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}c_{1}\textup{dist}(x_{i},X^{*})$ (46)
-$\leq(1+c_{1})\varepsilon+c_{1}\varepsilon\sum_{i=1}^{\infty}\eta^{i}=\Big{(}1+\frac{c_{1}}{1-\eta}\Big{)}\varepsilon\leq\frac{\delta}{2}.$
-Theorem 2.
-Suppose that Assumptions 1-3 are valid, and $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$.
-Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ for all $k$.
-If Assumption 4 is verified and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (45), then $\{\textup{dist}(x_{k},X^{*})\}$ converges superlinearly to zero.
-Moreover, the sequence $\{x_{k}\}$ converges to a solution $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
-Proof.
+Proof
+Same as that of Lemma 9. ∎
+Theorem 3
+Let $x^{*}\in X^{*}$, suppose that Assumptions 1, 2 and 4 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$, $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$.
+Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ for all $k$, and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (39).
+Then, $\{\textup{dist}(x_{k},X^{*})\}$ converges superlinearly to zero.
+Moreover, the sequence $\{x_{k}\}$ converges to some $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
+Proof
 From (38) we obtain
 $\dfrac{\textup{dist}(x_{k+1},X^{*})}{\textup{dist}(x_{k},X^{*})}\leq\dfrac{\tilde{C}}{\omega}\textup{dist}(x_{k},X^{*})^{r}$
-and since $\textup{dist}(x_{k},X^{*})^{r}\rightarrow 0$ (due to Lemmas 8 and 9) we have that $\{\textup{dist}(x_{k},X^{*})\}$ converges to zero superlinearly.
-Once $\textup{dist}(x_{k},X^{*})$ converges to zero and $x_{k}\in B(x^{*},\frac{\delta}{2})$ for every $k$, it remains to show that $\{x_{k}\}$ converges. From Lemma 5 and Lemma 8, we have
-$\|d_{k}\|\leq c_{1}\text{dist}(x_{k},X^{*})\leq c_{1}\eta^{k}\text{dist}(x_{0},X^{*})\leq c_{1}\varepsilon\eta^{k},$
-for every $k\geq 1$. Thus, for any positive integers $\ell$ and $q$, with $\ell\geq q$,
-$\|x_{\ell}-x_{q}\|\leq\sum_{i=q}^{\ell-1}\|d_{i}\|\leq\sum_{i=q}^{\infty}\|d_{i}\|\leq c_{1}\varepsilon\sum_{i=q}^{\infty}\eta^{i},$
-implying that $\{x_{k}\}\subset\mathbb{R}^{n}$ is a Cauchy sequence and, therefore, converges. ∎∎
-Remark 3.
-In particular, when $r=1$ in Assumption 4, $\textup{dist}(x_{k},X^{*})$ converges to zero quadratically.
+and since $\textup{dist}(x_{k},X^{*})^{r}\rightarrow 0$ (due to Lemmas 10 and 11) we have that $\{\textup{dist}(x_{k},X^{*})\}$ converges to zero superlinearly.
+Proof of convergence of the whole sequence $\{x_{k}\}$ to some $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$ follows the same lines as in the proof of Theorem 2. ∎
 4.2 Diminishing Rank
-Since the matrix rank is a lower semicontinuous function, for a given sequence $\{x_{k}\}$ converging to $x^{*}\in X^{*}$ it may happen that $\text{rank}J(x_{k})>\text{rank}J(x^{*})$, for every $k$ sufficiently large.
-We refer to this situation as “diminishing rank” scenario.
-In this case the convergence results will depend on the specific choice $\lambda_{k}=\|J(x_{k})^{T}F(x_{k})\|^{r^{\prime}}$, with $r^{\prime}\leq r$ from Assumption 4.
-To simplify the exposition, we consider $r^{\prime}=r$ in the results that follow.
-Lemma 10.
-Suppose that Assumptions 1-3 and 4 are valid, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
-$\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r}$ (47)
-with $\eta\in(0,1)$, then
+In this section, we will examine the scenario where the rank of $J(x_{k})$ decreases as $x_{k}$ approaches the set of stationary points $X^{*}$.
+In this case, the convergence analysis will depend not only on the upper bound for $\|J(x_{k})^{T}F(\bar{x}_{k})\|$ given in Assumption 3 (or Assumption 4) but also on the specific choices of the LM parameter $\lambda_{k}$.
+Lemma 12
+Let $x^{*}\in X^{*}$, suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$.
+Additionally, assume that Assumption 4 is verified with $r\in]0,1]$, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$.
+If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
+$\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r}$
+with $\eta\in(0,1)$, and $\hat{C}=L_{7}+(1+(1+c_{1})^{1+r})C$, $L_{7}=L_{4}c_{1}^{2}+L_{5}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}$, then
 $\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
-Proof.
-Note that from Assumption 4 and Lemma 6, for $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, we have $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$, with
-$c_{1}=\frac{1}{\gamma\omega^{r}}(L_{4}+C)+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}.$ (48)
-From Lemma 7, Assumption 4, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, (10), $\textup{dist}(x_{k},X^{*})\leq\delta/2<1/2$ and assuming that $\textup{dist}(x_{k},X^{*})<\varepsilon$, we have
-$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||+$ (49)
-$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||$
-$\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{1+r}$
+Proof
+Note that from Assumption 4 and Lemma 6(a), for $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, we have $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$, with
+$c_{1}=\frac{4}{\gamma\omega^{r}}(L_{4}+C)+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}.$ (40)
+From Lemma 7, Assumptions 2 and 4, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, $\textup{dist}(x_{k},X^{*})\leq\delta/2<1/2$ and assuming that $\textup{dist}(x_{k},X^{*})<\varepsilon$, we have
+$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{1+r}$
 $+C\|x_{k}-\bar{x}_{k}\|^{1+r}+C\|x_{k+1}-\bar{x}_{k}\|^{1+r}$
@@
 with $c_{1}$ as in (40) and $L_{7}=L_{4}c_{1}^{2}+L_{5}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}$.
-Therefore, for $\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r},$ we have $\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*})$. ∎∎
+Therefore, for $\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r},$ we have $\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*})$. ∎
 Using an appropriate $\varepsilon>0$, and Lemma 12 , the proofs of the next results are analogous to those of Lemma 11 and Theorem 3, respectively.
-Lemma 11.
+Lemma 13
 Suppose that the Assumptions of Lemma 12 are satisfied, and consider $\varepsilon$ as
@@
 If $x_{0}\in B(x^{*},\varepsilon)$, then $x_{k+1}\in B(x^{*},\frac{\delta}{2})$, and $\textup{dist}(x_{k},X^{*})\leq\varepsilon$ for every $k\in\mathbb{N}$.
-Theorem 3.
-Suppose that Assumptions 1-3 and 4 are valid.
-Let $\{x_{k}\}$ be generated by LMMSS with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, for all $k$.
-If $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ as in (50), then $\{dist(x_{k},X^{*})\}$ converges superlinearly to zero.
-Furthermore, the sequence $\{x_{k}\}$ converges to the solution $\bar{x}\in X^{*}\cap B(x^{*},\delta/2).$
-Remark 4.
+Theorem 4
+Let $x^{*}\in X^{*}$, suppose that Assumptions 1, 2 and 4 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$ and $r\in]0,1]$, and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$.
+Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ for all $k$, and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (41).
+Then, $\{\textup{dist}(x_{k},X^{*})\}$ converges superlinearly to zero.
+Moreover, the sequence $\{x_{k}\}$ converges to some $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
+Remark 4
 The main difference from Theorem 3 is that when Assumption 4 holds with $10<r<$ 1 we cannot choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ anymore; $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ is essential for the convergence proof in this case.
-This becomes clear from the proof of Lemma 6; see inequality (36).
+This becomes clear from the proof of Lemma 6(a); see inequality (31).
 In fact, we could choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r^{\prime}}$ with $r^{\prime}\leq r$ that all analysis remains valid with minor modifications.
+To close this section, we analyze local convergence of LMMSS when the Jacobian rank is not constant and Assumption 3 holds.
+Lemma 14
+Let $x^{*}\in X^{*}$, suppose that Assumptions 1 and 2 are valid in $B(x^{*},\delta)$ for some $\delta\in]0,1[$ and that the uniform completeness condition holds at $B(x^{*},\delta)\setminus X^{*}$.
+Also, assume that Assumption 3 is verified with
+$\sigma<\bar{\sigma}:=\min\left\{(\sigma^{*}_{\min})^{2},\frac{\omega}{(2+(1+\theta\|L\|^{2}))c_{1}}\right\},$
+and $\lambda_{k}$, $\theta$ and $\bar{\delta}$ are those in Lemma 6(b). If $x_{k},x_{k+1}\in B(x^{*},\bar{\delta}/2)$, and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
+$\varepsilon\leq\frac{\eta\omega-(2+(1+\theta\|L\|^{2})c_{1})\sigma}{L_{8}},$
+with $\eta\in\left]\frac{(2+(1+\theta\|L\|^{2})c_{1})\sigma}{\omega},1\right[$ and $L_{8}=L_{4}(c_{1}^{2}+\theta c_{1}\|L\|^{2})+L_{5}$, then
+$\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
+Proof
+From Assumption 3 and Lemma 6(b), for $\lambda_{k}$ as in (27), we have $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$, with
+$c_{1}=(1/\gamma)\max\{\kappa,(\sigma_{\min}^{*})^{2}/\theta\}+\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}/\sqrt{\gamma}.$ (42)
+From Lemma 7, Assumptions 2 and 3, $\textup{dist}(x_{k},X^{*})\leq\delta/2<1/2$ and assuming that $\textup{dist}(x_{k},X^{*})<\varepsilon$, we have
+$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||$
+$+\sigma\|x_{k}-\bar{x}_{k}\|+\sigma\|x_{k+1}-\bar{x}_{k}\|$
+$\leq(L_{4}(c_{1}^{2}+\theta c_{1}\|L\|^{2})+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+(2+(1+\theta\|L\|^{2})c_{1})\sigma\|x_{k}-\bar{x}_{k}\|$
+$\leq\Big{[}(L_{4}(c_{1}^{2}+\theta c_{1}\|L\|^{2})+L_{5})\varepsilon+(2+(1+\theta\|L\|^{2})c_{1})\sigma\Big{]}\|x_{k}-\bar{x}_{k}\|.$
+Now, for $\sigma<\frac{\omega}{(2+(1+\theta\|L\|^{2}))c_{1}}$, for any $\eta\in\left]\frac{(2+(1+\theta\|L\|^{2}))\sigma}{\omega},1\right[$, we have $\eta\omega-(2+(1+\theta\|L\|^{2})c_{1})\sigma>0$.
+Denoting $L_{8}=L_{4}(c_{1}^{2}+\theta c_{1}\|L\|^{2})+L_{5}$, for $\varepsilon\leq\frac{\eta\omega-(2+(1+\theta\|L\|^{2})c_{1})\sigma}{L_{8}}$, we obtain $\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*})$.
+∎
+Again, with the aid of Lemma 14, the proofs of the next results are analogous to those of Lemma 12 and Theorem 4, respectively.
+Lemma 15
+Suppose that the Assumptions of Lemma 14 are satisfied in $B(x^{*},\delta)$, consider $\bar{\delta}$ from Lemma 6(b) and
+$\varepsilon=\min\left\{\frac{\frac{\bar{\delta}}{2}}{1+\frac{c_{1}}{1-\eta}},\frac{\eta\omega-(2+(1+\theta\|L\|^{2})c_{1})\sigma}{L_{8}}\right\}.$ (43)
+If $x_{0}\in B(x^{*},\varepsilon)$, then $x_{k+1}\in B(x^{*},\frac{\bar{\delta}}{2})$, and $\textup{dist}(x_{k},X^{*})\leq\varepsilon$ for every $k\in\mathbb{N}$.
+Theorem 5
+Let $x^{*}\in X^{*}$, suppose that the assumptions of Lemma 14 hold in $B(x^{*},\delta)$.
+Let $\{x_{k}\}$ be generated by LMMSS with $\alpha_{k}=1$, $\lambda_{k}$ as in (27), for all $k$, $\theta,\bar{\delta}$ from Lemma 6(b).
+If $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ as in (43), then $\{dist(x_{k},X^{*})\}$ converges linearly to zero.
+Furthermore, the sequence $\{x_{k}\}$ converges to the solution $\bar{x}\in X^{*}\cap B(x^{*},\bar{\delta}/2)$.
+Remark 5
+In the results under Assumption 3 that led to Theorem 5, we considered that the Jacobian rank may change in a neighborhood of $X^{*}$.
+In this case, we emphasize that, in order to apply Lemma 6(b), $\lambda_{k}$ must follow the regime defined by (27), which implies $\sigma\leq\lambda_{k}\leq\theta L_{4}\delta+\theta\sigma$.
+Thus, in this setting, $\lambda_{k}$ must stays bounded away from zero, as opposed to the previous results in this work.
 5 Global convergence
-Now we consider Algorithm 1, a version of LMMSS globalized with line-search for non-zero residue nonlinear least-squares.
-It is worth to point out that Algorithm 1 differs from the algorithm proposed in [5] in the choice of the LM parameter $\lambda_{k}$ as well as the full step ($\alpha_{k}=1$) acceptance criterion.
-Input: $\nu,\zeta,\vartheta,\theta\in(0,1)$, $M>0$ $F$, $J$, $L$, and $x_{0}\in\Omega$.
-Set $\lambda_{0}=\left\lVert\nabla\phi({x}_{0})\right\rVert$, $k=0$.
+Now we consider Algorithm 1, a version of LMMSS globalized with line-search for nonzero residual nonlinear least-squares.
+It is worth to point out that Algorithm 1 differs from the algorithm proposed in Boos2024 in:
+  • the choice of the LM parameter $\lambda_{k}$,
+  • the full step ($\alpha_{k}=1$) acceptance criterion,
+  • a LMM direction as safeguard when LMMSS direction fails to satisfy sufficient descent conditions.
+Input: $\nu,\zeta,\vartheta,\xi\in]0,1[$, $M>0$, $F$, $J$, $L$, $\lambda_{0}>0$ and $x_{0}\in\mathbb{R}^{n}$.
+Set $k=0$.
   1. 1.
@@
   2. 2.
-    Compute $d_{k}$, a solution of $(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d=-J_{k}^{T}F_{k}$.
+    Choose $\lambda_{k}>0$ and compute $d_{k}$, a solution of $(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d=-J_{k}^{T}F_{k}$.
   3. 3.
@@
   4. 4.
-    If $\|d_{k}\|>M$ or $-\nabla\phi(x_{k})^{T}d_{k}<\theta\|\nabla\phi(x_{k})\|^{2}$, set
+    If $\|d_{k}\|>M$ or $-\nabla\phi(x_{k})^{T}d_{k}<\xi\|\nabla\phi(x_{k})\|^{2}$, set
     $d_{k}=-(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}J_{k}^{T}F_{k}.$
   5. 5.
-    Choose $m$ as the smallest non-negative integer such that
+    Choose $m$ as the smallest nonnegative integer such that
     $\phi(x_{k}+\zeta^{m}{d}_{k})-\phi({x}_{k})\leq\nu\zeta^{m}\nabla\phi({x}_{k})^{T}{d}_{k}$
@@
   6. 6.
-    Update ${x}_{k+1}={x}_{k}+\alpha_{k}{d}_{k}$, $\lambda_{k+1}=\left\lVert\nabla\phi(x_{k+1})\right\rVert$, $k\leftarrow k+1$,
+    Update ${x}_{k+1}={x}_{k}+\alpha_{k}{d}_{k}$, $k\leftarrow k+1$,
 and go to Step 1.
-Algorithm 1 LMMSS with line-search
-Initially, we will demonstrate that the sequence of directions $d_{k}$ generated by Algorithm 1 is gradient-related (see [4, Sec.
-1.2]).
+Algorithm 1 Globalized LMMSS
+As stated, Algorithm 1 leaves open the choice of the sequence of positive LM parameters $\{\lambda_{k}\}$.
+Except for the very last theorem of this section, all the results that follow hold for the choices $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r},r\in]0,1]$ or $\lambda_{k}\in[\underline{\lambda},\bar{\lambda}]$, $0<\underline{\lambda}\leq\bar{\lambda}<+\infty$.
+[…]
+Initially, we will demonstrate that the direction sequence $\{d_{k}\}$ generated by Algorithm 1 is gradient-related to $\{x_{k}\}$ (see (Bertsekas2016, , Eq(1.14))).
 Subsequently, we will establish that any limit point of the sequence produced by this algorithm is a stationary point for (1), regardless of the initial point.
-In this section, we assume that Assumption 4 is satisfied with $r=1$ to simplify notation and proofs.
-We recall the definition of gradient-related directions from [4, Sec. 1.2].
-Definition 1 (Gradient-related).
-Let $\{x_{k}\}$ and $\{d_{k}\}$ be sequences in $\mathbb{R}^{n}$.
-The sequence $\{d_{k}\}$ is said to be gradient-related to $\{x_{k}\}$ if, for each subsequence $\{x_{k}\}_{k\in\mathcal{K}}$ (with $\mathcal{K}\subseteq\mathbb{N}$) converging to a non-stationary point of $\phi$, the corresponding subsequence $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded and satisfies
-$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}<0.$ (51)
-Proposition 1.
+We recall the definition of gradient-related directions from (Bertsekas2016, , Eq(1.14)).
+Definition 2 (Gradient-related)
+Let $\{x_{k}\}$ and $\{d_{k}\}$ be sequences in $\mathbb{R}^{n}$ and $\phi:\mathbb{R}^{n}\to\mathbb{R}$ a continuously differentiable function.
+The sequence $\{d_{k}\}$ is said to be gradient-related to $\{x_{k}\}$ if, for each subsequence $\{x_{k}\}_{k\in\mathcal{K}}$ (with $\mathcal{K}\subseteq\mathbb{N}$) converging to a nonstationary point of $\phi$, the corresponding subsequence $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded and satisfies
+$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}<0.$
+Remark 6
+Let $\{d_{k}\}$, $\{x_{k}\}$ be sequences generated by an iterative algorithm $x_{k+1}=x_{k}+\alpha_{k}d_{k}$, for minimizing $\phi(x)$, where $\alpha_{k}$ fulfills Armijo condition.
+It is not hard to prove that if $\{d_{k}\}$, $\{x_{k}\}$ satisfy
+$\|d_{k}\|\leq\bar{M}\quad\text{and}\quad\nabla\phi(x_{k})^{T}d_{k}\leq-\bar{\xi}\|\nabla\phi(x_{k})\|^{p_{1}},$
+for $\bar{M}>0$, $p_{1}>0$ and $\bar{\xi}\in]0,1[$, then $\{d_{k}\}$ is gradient-related.
+Proposition 1
 Let $\{d_{k}\}$, $\{x_{k}\}$ be sequences generated by Algorithm 1 starting from $x_{0}\in\mathbb{R}^{n}$.
-Suppose that Assumption 1 holds in the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$.
+Suppose that condition (4) holds in the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$, and that either $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ or $0<\underline{\lambda}\leq\lambda_{k}\leq\bar{\lambda}$.
 Then, $\{d_{k}\}$ is gradient-related.
-Proof.
-Suppose that Algorithm 1 generates an infinite sequence $\{x_{k}\}$, i.e., $\|\nabla\phi(x_{k})\|=\|J_{k}^{T}F_{k}\|>0$, for each $k$, then, consequently, $\left\lVert F(x_{k})\right\rVert>0$, and let $\{x_{k}\}_{k\in\mathcal{K}}$, $\mathcal{K}\subseteq\mathbb{N}$, be a subsequence converging to $x_{\infty}$, a non-stationary point for $\phi$, i.e., $\nabla\phi(x_{\infty})\neq 0$.
+Proof
+Suppose that Algorithm 1 generates an infinite sequence $\{x_{k}\}$, i.e., $\|\nabla\phi(x_{k})\|=\|J_{k}^{T}F_{k}\|>0$, for each $k$.
+Then, $\left\lVert F(x_{k})\right\rVert>0$ in this sequence.
+Let $\{x_{k}\}_{k\in\mathcal{K}}$, $\mathcal{K}\subseteq\mathbb{N}$, be a subsequence converging to $\hat{x}$, a nonstationary point for $\phi$, i.e., $\nabla\phi(\hat{x})\neq 0$.
 From now on, consider $k\in\mathcal{K}$.
@@
 $\nabla\phi(x_{k})^{T}d_{k}$ $=-d_{k}^{T}(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}$
-$=-d_{k}^{T}X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}d_{k}$
-$:=-d_{k}^{T}X_{k}^{-T}\Lambda_{k}X_{k}^{-1}d_{k}.$
-By Ostrowski’s inertia law, we have
+$=-d_{k}^{T}X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}d_{k}$ (45)
+$:=-d_{k}^{T}X_{k}^{-T}\Lambda_{k}X_{k}^{-1}d_{k}.$ (46)
+By Ostrowski’s inertia law (ostrowski, , Theorem 1), we have
 $\lambda_{n}(X_{k}^{-T}\Lambda_{k}X_{k}^{-1})=\sigma_{n}^{2}(X_{k}^{-1})\lambda_{n}(\Lambda_{k}),$
@@
 Then, from Lemma 3 it follows that
-$\lambda_{n}(X_{k}^{-T}\Lambda_{k}X_{k}^{-1})\geq\gamma\lambda_{n}(\Lambda_{k})\geq\gamma\min\{1,\lambda_{k}\},$
+$\lambda_{n}(X_{k}^{-T}\Lambda_{k}X_{k}^{-1})\geq\gamma\lambda_{n}(\Lambda_{k})\geq\gamma\min\{1,\lambda_{k}\},$ (47)
 where, in the last inequality, we use the fact that $\sigma_{i,k}^{2}+\mu_{i,k}^{2}=1$.
-Since $x_{\infty}$ is non-stationary, and $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert$, there exists a $\underline{\lambda}>0$ such that for $k\in{\cal K}$, $\lambda_{k}\geq\underline{\lambda}$.
-Thus, we obtain
+If $\lambda_{k}$ is chosen such that $\lambda_{k}\geq\underline{\lambda}>0$, from (46) and (47) we obtain
 $\nabla\phi(x_{k})^{T}d_{k}\leq-\gamma\min\{1,\underline{\lambda}\}\|d_{k}\|^{2}.$ (48)
+If $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, (48) also holds, because $\left\lVert J_{k}^{T}F_{k}\right\rVert>0$, for $k\in{\cal K}$, and $\hat{x}$ is nonstationary, which implies the existence of $\underline{\lambda}>0$ such that for $k\in{\cal K}$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|\geq\underline{\lambda}$.
 On the other hand,
@@
 $\leq\left(\|J_{k}\|^{2}+\lambda_{k}\|L\|^{2}\right)\|d_{k}\|.$
-By the continuity of $F(x)$ and $J(x)$ and the convergence of the subsequence $\{x_{k}\}_{\mathcal{K}}$, there exist $L_{J}>0$ and $M_{0}>0$ such that $\|J(x_{k})^{T}F(x_{k})\|\leq M_{0}$ and $\left\lVert J(x_{k})\right\rVert\leq L_{J}$ for each $k$.
-Thus, since $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert$, we have
-$\|\nabla\phi(x_{k})\|\leq\left(L_{J}^{2}+M_{0}\|L\|^{2}\right)\|d_{k}\|=:M_{1}\|d_{k}\|,\quad\forall k\in{\cal K}.$ (53)
+By the continuity of $F(x)$ and $J(x)$ and the convergence of the subsequence $\{x_{k}\}_{\mathcal{K}}$, there exist $L_{J}>0$ and $M_{0}>0$ such that $\|J(x_{k})^{T}F(x_{k})\|\leq M_{0}$ and $\left\lVert J(x_{k})\right\rVert\leq L_{J}$ for each $k\in{\cal K}$.
+Thus, if either $0<\underline{\lambda}\leq\lambda_{k}\leq\bar{\lambda}$ or $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}\leq M_{0}^{r}=:\bar{\lambda}$, we have
+$\|\nabla\phi(x_{k})\|\leq\left(L_{J}^{2}+\bar{\lambda}\|L\|^{2}\right)\|d_{k}\|=:M_{1}\|d_{k}\|,\quad\forall k\in{\cal K}.$ (49)
 From (49) and (48), we obtain
@@
 which implies that
-$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}\leq-\frac{\gamma\min\{1,\underline{\lambda}\}}{M_{1}^{2}}\|\nabla\phi(x_{\infty})\|^{2}<0.$
-Now we just need to show that $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded. Recall from (19) that
-$\|d_{k}\|\leq\|X_{k}\|^{2}\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|,\quad\forall k\geq 0,$ (54)
+$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}\leq-\frac{\gamma\min\{1,\underline{\lambda}\}}{M_{1}^{2}}\|\nabla\phi(\hat{x})\|^{2}<0.$
+Now we just need to show that $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded.
+Recall from (14) that
+$\|d_{k}\|\leq\|X_{k}\|^{2}\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|,\quad\forall k\geq 0,$
 with $\Gamma_{k}=(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}$.
-From Lemma 3, we have $\|X_{k}\|\leq 1/\sqrt{\gamma}$. Now let us analyze the term $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|$. From Lemma 4, it follows that
-$\|\Gamma_{k}\|\leq\dfrac{1}{\lambda_{k}},\quad\text{if}\quad 0<\lambda_{k}<1,\quad\text{and}\quad\|\Gamma_{k}\|\leq 1,\quad\text{if}\quad\lambda_{k}\geq 1.$ (55)
+From Lemma 3, we have $\|X_{k}\|\leq 1/\sqrt{\gamma}$. Now let us analyze the term $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|$.
+From Lemma 4, it follows that
+$\|\Gamma_{k}\|\leq\dfrac{1}{\lambda_{k}},\quad\text{if}\quad 0<\lambda_{k}<1,\quad\text{and}\quad\|\Gamma_{k}\|\leq 1,\quad\text{if}\quad\lambda_{k}\geq 1.$
 We separate into two cases to analyze such upper bound.
   • (a)
-    If $10<\lambda_{k}<$ 1, then $\max\{\|\Gamma_{k}\|,1\}\leq\dfrac{1}{\lambda_{k}}$. Since $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, we have
-    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\dfrac{1}{\lambda_{k}}\|J_{k}^{T}F_{k}\|=\dfrac{1}{\|J_{k}^{T}F_{k}\|}\|J_{k}^{T}F_{k}\|=1.$
+    If $10<\lambda_{k}<$ 1, then
+    $\max\{\|\Gamma_{k}\|,1\}\leq\dfrac{1}{\lambda_{k}}.$
+    If $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, we have
+    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\dfrac{1}{\lambda_{k}}\|J_{k}^{T}F_{k}\|=\dfrac{1}{\|J_{k}^{T}F_{k}\|^{r}}\|J_{k}^{T}F_{k}\|=\|J_{k}^{T}F_{k}\|^{1-r}.$
+    If $\lambda_{k}\geq\underline{\lambda}>0$, we obtain
+    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\dfrac{1}{\lambda_{k}}\|J_{k}^{T}F_{k}\|=\dfrac{1}{\underline{\lambda}}\|J_{k}^{T}F_{k}\|.$
+    In both cases the right hand side is bounded by $\max\left\{M_{0}^{1-r},M_{0}/\underline{\lambda}\right\}$.
   • (b)
-    If $\lambda_{k}\geq 1$, then $\max\{\|\Gamma_{k}\|,1\}\leq 1$. Thus, $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq M_{0}$.
+    If $\lambda_{k}\geq 1$, then $\max\{\|\Gamma_{k}\|,1\}\leq 1$. Thus,
+    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\|J_{k}^{T}F_{k}\|\leq M_{0},\quad\forall k\in{\cal K}.$
 Therefore, from (a) and (b), we conclude that
-$\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\max\{1,M_{0}\},\quad\forall k\in{\cal K}.$
+$\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\max\{M_{0}^{1-r},M_{0}/\underline{\lambda},M_{0}\}=:\bar{M}_{0},\quad\forall k\in{\cal K}.$
 and
-$\|d_{k}\|\leq\frac{1}{\gamma}\max\{1,M_{0}\}=:M_{2},\quad\forall k\in{\cal K}.$ (56)
-Inequalities (52) and (56) show that for $k\in{\cal K}$, $d_{k}$ is the one determined in Step 2 of Algorithm 1 because it passes the seize of Step 4 for $M=M_{2}$ and $\theta=\gamma\min\{1,\underline{\lambda}\}/M_{1}^{2}$.
-Thus, the proof is complete.
-∎∎
-Remark 5.
-The above proposition shows that LMMSS directions generated by Algorithm 1 are gradient-related as long as the completeness condition (Assumption 1) holds at every iterate – which also ensures that the solution of (2) is unique.
-Nevertheless, apart from special cases, e.g when the Jacobian $J(x)$ is full-rank everywhere, we cannot ensure that Assumption 1 holds at every $x_{k}$ generated by Algorithm 1.
-Hence, we safeguard the algorithm by using the classic LMM direction whenever the LMMSS direction fails to fulfill sufficient conditions for gradient-related property, namely, $\|d_{k}\|\leq M$, for $M>0$ and $-\nabla\phi(x_{k})^{T}d_{k}\geq\theta\|\nabla\phi(x_{k})\|^{p_{1}}$, for $\theta\in(0,1)$ and $p_{1}>0$.
+$\|d_{k}\|\leq\frac{1}{\gamma}\max\{\bar{M}_{0},M_{0}\}=:M_{2},\quad\forall k\in{\cal K}.$
+Thus, the proof is complete. ∎
+The above proposition shows that directions $d_{k}$ generated at Step 2 of Algorithm 1 are gradient-related, for both choices $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, or $0<\underline{\lambda}\leq\lambda_{k}\leq\bar{\lambda}$, as long as the completeness condition (4) holds at every iterate.
+Nevertheless, apart from special cases, e.g when the Jacobian $J(x)$ is full-rank everywhere, we cannot ensure that such condition holds at every $x_{k}$ generated by Algorithm 1 starting at an arbitrary $x_{0}\in\mathbb{R}^{n}$.
+Hence, we safeguard the algorithm by using the classic LMM direction whenever the LMMSS direction fails to fulfill sufficient conditions for gradient-related property, namely, those of Remark 6.
 If such conditions are not satisfied, we take $d_{k}=-(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}J_{k}^{T}F_{k}$ which yields gradient-related directions. In fact, from
 $(J_{k}^{T}J_{k}+\lambda_{k}I)d_{k}=-\nabla\phi(x_{k}),$
-we obtain $-\nabla\phi(x_{k})^{T}d_{k}=\|J_{k}d_{k}\|^{2}+\lambda_{k}\|d_{k}\|^{2}\geq\lambda_{k}\|d_{k}\|^{2},$ and since
+we obtain
+$-\nabla\phi(x_{k})^{T}d_{k}=\|J_{k}d_{k}\|^{2}+\lambda_{k}\|d_{k}\|^{2}\geq\lambda_{k}\|d_{k}\|^{2},$
+and since
 $\|\nabla\phi(x_{k})\|=\|(J_{k}^{T}J_{k}+\lambda_{k}I)d_{k}\|\leq\|J_{k}^{T}J_{k}+\lambda_{k}I\|\|d_{k}\|\leq M_{3}\|d_{k}\|,$
-where $M_{3}=L_{J}^{2}+M_{0}$. We conclude that
-$-\nabla\phi(x_{k})^{T}d_{k}\geq\lambda_{k}\|d_{k}\|^{2}\geq\frac{\lambda_{k}}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2}=\frac{1}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{3}$
-if we take $\lambda_{k}=\|\nabla\phi(x_{k})\|$. Moreover,
-$\|d_{k}\|=\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\nabla\phi(x_{k})\|$ $\leq\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\|\|\nabla\phi(x_{k})\|\leq 1,$
-which shows that our “safeguard directions” are also gradient-related.
-Now that it is proven that Algorithm 1 generates a sequence of directions $\{d_{k}\}$ that is gradient-related, using [4, Proposition 1.2.1], global convergence can be established.
-Theorem 4.
-Let $\{x_{k}\}$ be a sequence generated by Algorithm 1. Then, every limit point $\hat{x}$ of $\{x_{k}\}$ is such that $\nabla\phi(\hat{x})=0$.
-Proof.
+where $M_{3}=L_{J}^{2}+\bar{\lambda}$. We conclude that
+$-\nabla\phi(x_{k})^{T}d_{k}\geq\lambda_{k}\|d_{k}\|^{2}\geq\frac{\lambda_{k}}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2}=\frac{1}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2+r}$
+if we take $\lambda_{k}=\|\nabla\phi(x_{k})\|^{r}$, or $-\nabla\phi(x_{k})^{T}d_{k}\geq\frac{\underline{\lambda}}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2}$, if $\lambda_{k}\geq\underline{\lambda}>0$.
+Moreover,
+$\|d_{k}\|=\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\nabla\phi(x_{k})\|$ $\leq\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\|\|\nabla\phi(x_{k})\|$
+$\leq\frac{1}{\lambda_{k}}\|\nabla\phi(x_{k})\|=\|\nabla\phi(x_{k})\|^{1-r},$
+if $\lambda_{k}=\|\nabla\phi(x_{k})\|^{r}$, or $\|d_{k}\|\leq\|\nabla\phi(x_{k})\|/\underline{\lambda}$ if $\lambda_{k}\geq\underline{\lambda}>0$.
+In both cases the right rand sides are bounded in the convergent subsequence $\{x_{k}\}_{k\in{\cal K}}$.
+Now that it is proven that Algorithm 1 generates a sequence of directions $\{d_{k}\}$ that is gradient-related, using (Bertsekas2016, , Proposition 1.2.1), global convergence can be established.
+Theorem 6
+Given $x_{0}\in\mathbb{R}^{n}$, let $\{x_{k}\}$ be a sequence generated by Algorithm 1 with either $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r},r\in]0,1[$ or $0<\underline{\lambda}\leq\lambda_{k}\leq\bar{\lambda}$.
+Then, every limit point $\hat{x}$ of $\{x_{k}\}$ is such that $\nabla\phi(\hat{x})=0$.
+Proof
 Let $K_{1}=\{k\in\mathbb{N}\mid\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|\leq\vartheta\|J(x_{k})^{T}F(x_{k})\|\}$.
@@
 Otherwise, if $K_{1}$ is finite, let us assume, without loss of generality, that $\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|>\vartheta\|J(x_{k})^{T}F(x_{k})\|$, for each $k$, such that the step size is chosen to satisfy the Armijo condition.
-Since, by Proposition 1 and Remark 5, the directions of Algorithm 1 are gradient-related, it follows from [4, Proposition 1.2.1] that any limit point $\hat{x}$ of $\{x_{k}\}$ is a stationary point of $\phi(x)$.
-∎∎
+Since the directions of Algorithm 1 are gradient-related (due to Proposition 1 and discussion afterwards), it follows from (Bertsekas2016, , Proposition 1.2.1) that any limit point $\hat{x}$ of $\{x_{k}\}$ is a stationary point of $\phi(x)$.
+∎
 The next result establishes the connection between global and local convergence by showing that, under certain conditions, $\alpha_{k}=1$ for all $k$ sufficiently large and then Algorithm 1 becomes the “pure” LMMSS and the results of Section 4 are applicable.
-Theorem 5.
-Let $\{x_{k}\}$ be generated by Algorithm 1, using $x_{0}\in\mathbb{R}^{n}$ as the initial point.
-Suppose that Assumptions 1-3 and Assumption 4, with $r=1$, hold at the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$.
+This is the only result in this section which requires Assumption 4.
+Theorem 7
+Let $\{x_{k}\}$ be generated by Algorithm 1, using $x_{0}\in\mathbb{R}^{n}$ as the initial point and $\lambda_{k}=\|J_{k}^{T}F_{k}\|$.
+Suppose that Assumption 2, Assumption 4 with $r=1$, and condition (4) hold at the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$.
 Moreover, assume that ${\cal L}_{0}$ is compact.
-Then, for every $k$ sufficiently large $\alpha_{k}=1$ and the sequence $\{\textup{dist}(x_{k},X^{*})\}$ converges to 0 quadratically.
-Proof.
+Then, for every $k$ sufficiently large, $\alpha_{k}=1$, and the sequence $\{\textup{dist}(x_{k},X^{*})\}$ converges quadratically to zero.
+Proof
 Note that by the definition of Algorithm 1 $\{\phi(x_{k})\}$ is decreasing.
 Starting from $x_{0}$, we conclude that the sequence $\{x_{k}\}$ remains in ${\cal L}_{0}$, a compact set.
-Therefore, there exists a convergent subsequence, as per Theorem 4, where every limit point of this subsequence is stationary.
-By removing this convergent subsequence, we can repeat this process successively for the remaining sequence, which remains in a compact set, thereby concluding that $\{x_{k}\}$ is the disjoint and enumerable union of convergent subsequences, each of which has a stationary point as limit.
-Thus, for every $k$ sufficiently large, the norm of the gradient at $x_{k}$ will be sufficiently small. In other words, there is a positive integer $k_{0}$ such that
+By Theorem 6, every limit point $\hat{x}$ of $\{x_{k}\}$ is stationary, thus for $k$ sufficiently large, the norm of the gradient at $x_{k}$ will be sufficiently small.
+In other words, there is a positive integer $k_{0}$ such that
 $\|J(x_{k})^{T}F(x_{k})\|\leq\frac{\omega^{3}\vartheta}{\tilde{C}L_{3}},\quad\text{for all}\quad k\geq k_{0}.$ (50)
@@
 $=L_{3}\textup{dist}(x_{k+1},X^{*})\leq\frac{\tilde{C}}{\omega}L_{3}\textup{dist}(x_{k},X^{*})^{2}$
-$\leq\frac{\tilde{C}L_{3}\|J(x_{k})^{T}F(x_{k})\|}{\omega^{3}}\|J(x_{k})^{T}F(x_{k})\|$ (58)
+$\leq\frac{\tilde{C}L_{3}\|J(x_{k})^{T}F(x_{k})\|}{\omega^{3}}\|J(x_{k})^{T}F(x_{k})\|$
 $\leq\vartheta\|J(x_{k})^{T}F(x_{k})\|,$
 proving that $\alpha_{k}=1$ for each $k\geq k_{0}$.
-To obtain the quadratic convergence of $\{\textup{dist}(x_{k},X^{*})\}_{k\geq k_{0}}$ to 0, one may apply Theorem 3 when we have a diminishing rank or from Theorem 2, when $rank(J(x))=rank(J(x^{*}))\geq 1$.
-∎∎
-Remark 6.
-We observe that in the case of zero residue a previous result [13, Theorem 3.1] assumed that the limit point $x^{*}$ is such that $F(x^{*})=0$.
-Also, in the case of unconstrained optimization, in order to prove a result similar to Theorem 5, the required assumption is that the Hessian at $x^{*}$ is positive definite; however, such assumption would imply that $x^{*}$ is an isolated stationary point.
-As we want to address the case of non-zero residue and possibly non-isolated stationary points, we considered the boundness of the level set instead of these other two conditions.
-Remark 7.
+To obtain the quadratic convergence of $\{\textup{dist}(x_{k},X^{*})\}_{k\geq k_{0}}$ to 0, one may now apply Theorem 3 (or Theorem 4) with $x^{*}=\hat{x}$ and $x_{0}=x_{k_{0}}$.
+∎
+Remark 7
+We observe that, in the case of zero residual, a previous result (Yamashita, , Theorem 3.1) assumed that the limit point $x^{*}$ is such that $F(x^{*})=0$.
+Also, in the case of unconstrained optimization, in order to prove a result similar to Theorem 7 the required assumption is that the Hessian at $x^{*}$ is positive definite; however, such assumption would imply that $x^{*}$ is an isolated stationary point.
+As we want to address the case of nonzero residual and possibly nonisolated stationary points, we considered the boundness of the level set instead of these other two conditions.
+Remark 8
 In the scenario of diminishing rank, if Assumption 4 is satisfied with $10<r<$ 1, it is necessary to adjust the parameter $\lambda_{k}$ of Algorithm 1 to $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$.
-Furthermore, when this hypothesis holds, the above results follow with minor changes in some constants, and we can show there exists $k_{0}$ such that $\alpha_{k}=1,\forall k\geq k_{0}$ and that $\textup{dist}(x_{k},X^{*})$ goes to zero superlinearly.
+Furthermore, when this hypothesis holds, Theorem 7 follows with minor changes in some constants, and we can show there exists $k_{0}$ such that $\alpha_{k}=1,\forall k\geq k_{0}$ and that $\textup{dist}(x_{k},X^{*})$ goes to zero superlinearly.
 6 Illustrative examples
-The purpose of this section is to illustrate our convergence results by presenting some examples of non-zero residue nonlinear least-squares problems.
+The purpose of this section is to illustrate our convergence results by presenting some examples of nonzero residual nonlinear least-squares problems.
 In this section we use the shorthand notation (A1) – (A4) for Assumptions 14, respectively.
-Example 1. Let us start with an example from [2, Example 5.1] with non-isolated minimizers (stationary-points). Consider the residual function
+The examples came from reference (behling2019local, , Section 5), where (A1)–(A4) were already demonstrated/verified and such verification will not be repeated here.
+Example 1. (Constant rank and (A4)) Let us start with an example from (behling2019local, , Example 5.1) with nonisolated minimizers (stationary-points). Consider the residual function
 $F(x)=(x_{1}^{2}+x_{2}^{2}-1,x_{1}^{2}+x_{2}^{2}-9)^{T}.$
@@
 $J(x)=\left[\begin{array}[]{cc}2x_{1}&2x_{2}\\ 2x_{1}&2x_{2}\end{array}\right]$
-has rank $1$ everywhere except for the origin. Moreover, as discussed in [2], Assumptions (A2), (A3) and (A4) with $r=1$ hold as well.
-By considering LMMSS with the scaling matrix $L=[-1\quad 1]$, Assumption (A1) holds everywhere except for the line $x_{2}=-x_{1}$.
-Hence, we are under conditions to apply the results of Section 4.1.
-As expected from Theorem 2, in Table 1, we can observe the local quadratic convergence of $\text{dist}(x_{k},X^{*})$ to zero starting from two distinct starting points in a neighborhood of $x^{*}=(0,\sqrt{5})^{T}\in X^{*}$.
+has rank $1$ everywhere except for the origin.
+Moreover, as discussed in behling2019local , (A1), (A2) and (A4) with $r=1$ hold as well.
+It is worth to point out that (A4) holds with $C=0$ because in this example $\|J(x)^{T}F(z)\|=0$.
+By considering LMMSS with the scaling matrix $L=[-1\quad 1]$, condition (4) holds everywhere except for the line $x_{2}=-x_{1}$.
+Hence, if we consider an initial point in a neighborhood of $x^{*}=(0,\sqrt{5})^{T}\in X^{*}$, we are under conditions to apply the results of Section 4.1.
+As expected from Theorem 3, in Table 1, we can observe the local quadratic convergence of $\text{dist}(x_{k},X^{*})$ to zero, starting LMMSS from two distinct starting points in a neighborhood of $(0,\sqrt{5})^{T}$.
 Table 1: Stopping criterion $\|J_{k}^{T}F_{k}\|<10^{-8}$ with $x_{0}=(0,\sqrt{5}+0.03)^{T}$ as starting point (left), and $x_{0}=(0.01,\sqrt{5}-0.01)^{T}$ (right).
@@
 Figure 2a depicts the trajectory of LMM ($L^{T}L=I$) and the one of LMMSS (with $L=[-1\quad 1]$) both starting from $x_{0}=(2,4)^{T}$.
-As we can see from the figure, while LMM dives perpendicularly to the set $X^{*}$ (observe the level curves in the figure), LMMSS trajectory is parallel to the null space of $L$, and they end up at different stationary points.
-As exposed in [5] for parameter identification problems, an advantage of using Algorithm 1 is that in a given problem with non-isolated solutions, we can choose $L$ that enforces a specific property.
-In this geometric example, by using $L=[-1,1]$, we aim for the method to prioritize stationary points where $\|L(x-x_{0})\|=|x_{2}-x_{1}-2|$ is as small as possible.
-In [5], the authors employed $L$ as discrete versions of derivative operators in order to promote smoothness in the obtained solutions for parameter identification problems in heat conduction.
-Moreover, Figure 1b enforces this aspect by showing that the presence of a well-thought choice of $L$ may conduce the iterations faster and in a more accurate manner than LMM, while it stimulates a desired property.
+As we can see from the figure, while LMM dives perpendicularly to the set $X^{*}$ (observe the level curves in the figure), LMMSS trajectory seems parallel to the null space of $L$, and they end up at different stationary points but with the same residual norm.
+Figure 2b plots the gradient norm through the iterations of LMM and LMMSS.
 (a) Iterations and level curves of the objective function.
@@
 Figure 2: Plot of iterations for LMM and LMMSS for Example 1, with starting point $(2,4)$.
-The last aspect we shall discuss with this example is about the completeness condition and the global convergence mechanisms of Algorithm 1.
-As mentioned before, (A1) holds everywhere except for the line $x_{2}=-x_{1}$.
+The last aspect we shall discuss with this example is about the completeness condition (4) and the global convergence mechanisms of Algorithm 1.
+As mentioned before, condition  (4) holds everywhere except for the line $x_{2}=-x_{1}$.
 LMMSS also tries to move parallel to the null space of $L$.
@@
 The displacements $x_{k}-x_{0}$ are in ${\cal N}(L)$, but the hyperplane $x_{0}+{\cal N}(L)$ does not intersect $X^{*}$.
-As a result, LMMSS minimizes $\phi(x)$ in $x_{0}+{\cal N}(L)$ approximating a point $\hat{x}$ where $\nabla\phi(\hat{x})\perp{\cal N}(L)$ so that the directions $d_{k}$ become orthogonal to $-\nabla\phi(x_{k})$ loosing the descent property and the algorithm crashes not satisfying the Armijo condition.
-Notice that $\hat{x}$ is such that $\hat{x}_{2}=-\hat{x}_{1}$ and (A1) fails.
-On the other hand, we use Algorithm 1 with the safeguard step, Step 4 corrects this behavior and when $x_{k}$ approaches $\hat{x}$ the classic LMM direction is employed instead; see Figure 2b.
+As a result, LMMSS minimizes $\phi(x)$ in $x_{0}+{\cal N}(L)$, approximating a point $\hat{x}$ where $\nabla\phi(\hat{x})\perp{\cal N}(L)$ so that the directions $d_{k}$ become orthogonal to $-\nabla\phi(x_{k})$ loosing the descent property and the algorithm crashes not satisfying the Armijo condition.
+Notice that $\hat{x}$ is such that $\hat{x}_{2}=-\hat{x}_{1}$ and condition (4) fails.
+On the other hand, if we use Algorithm 1 with the safeguard step, Step 4 corrects this behaviour and, when $x_{k}$ approaches $\hat{x}$, the classic LMM direction is employed instead; see Figure 3b.
 This allows the algorithm to move on and reach $X^{*}$.
@@
 Figure 3: Plot of iterations for LMM and LMMSS for Example 1, with starting point $(-1,3)$, and level curves of the objective function.
-Example 2. Now we consider an example to illustrate the “diminishing rank” scenario. The residual function is given by
+Example 2. (Diminishing rank and (A4)) Now we consider an example to illustrate the “diminishing rank” scenario. The residual function is given by
 $F(x)=(x_{1}^{3}-x_{1}x_{2}+1,x_{1}^{3}+x_{1}x_{2}+1),$
-which leads to $\phi(x)$ having an isolated global minimizer at $(-1,0)^{T}$ and a non-isolated set of local minimizers $\{x\in\mathbb{R}^{2}\mid x_{1}=0\}$ that we shall use as $X^{*}$ and, in this case, $\textup{dist}(x,X^{*})=|x_{1}|$.
-This example was also analyzed in [2] where it was verified that assumptions (A2), (A3) and (A4) with $r=1$ hold in a neighborhood of $x^{*}=(0,2)^{T}$.
-For assumption (A1), we point out that the null space of
-$J(x)=\left[\begin{array}[]{cc}3x_{1}^{2}-x_{2}&-x_{1}\\ 3x_{1}^{2}+x_{2}&x_{1}\end{array}\right]$
-is $\{0\}$ wherever $x_{1}\neq 0$ and it is $\text{span}\{(0,1)^{T}\}$ for $x\in X^{*}\setminus\{0\}$. Hence, it is clear that, regardless the choice of $L$, (A1) holds everywhere except for $X^{*}$.
-In Table 2 and Figure 3 we report the behavior of LMM and LMMSS with $L=[-1\quad 1]$ starting in the neighborhood of $x^{*}=(0,2)^{T}$.
-From Table 2 the quadratic convergence rate is apparent for both methods.
-In Figure 3 we highlight the different limit points: while the classic LMM moves perpendicularly to $X^{*}$, LMMSS moves parallel to ${\cal N}(L)$, leading in this case to a stationary point with smaller $x_{2}$ coordinate.
+which leads to $\phi(x)$ having an isolated global minimizer at $(-1,0)^{T}$ and a nonisolated set of local minimizers $\{x\in\mathbb{R}^{2}\mid x_{1}=0\}$ that we shall use as $X^{*}$ and, in this case, $\textup{dist}(x,X^{*})=|x_{1}|$.
+This example was also analyzed in (behling2019local, , Example 5.2) where it was verified that assumptions (A1), (A2) and (A4) with $r=1$ hold in a neighborhood of $x^{*}=(0,2)^{T}$.
+Differently from the previous example where $J(x)^{T}F(z)=0$, here (A4) holds because $\|J(x)^{T}F(z)\|=6x_{1}^{2}=6\textup{dist}(x,X^{*})^{2}$.
+For condition (4), we point out that the null space of
+$J(x)=\left[\begin{array}[]{cc}3x_{1}^{2}-x_{2}&\quad-x_{1}\\ 3x_{1}^{2}+x_{2}&\quad x_{1}\end{array}\right]$
+is $\{0\}$ wherever $x_{1}\neq 0$ and it is $\text{span}\{(0,1)^{T}\}$ for $x\in X^{*}\setminus\{0\}$.
+Hence, it is clear that, regardless the choice of $L$, condition (4) holds everywhere except for $X^{*}$.
+In Table 2 and Figure 4 we report the behaviour of LMM and LMMSS, with $L=[-1\quad 1]$, starting in the neighbourhood of $x^{*}=(0,2)^{T}$.
+From Table 2, the quadratic convergence rate is apparent for both methods.
+In Figure 4 we highlight the different limit points: while the classic LMM moves almost perpendicularly to $X^{*}$, LMMSS moves parallel to ${\cal N}(L)$, leading in this case to a stationary point with smaller $x_{2}$ coordinate.
 Figure 4: Plot of iterations for LMM and LMMSS for Example 2, with starting point $(0.8,2.1)^{T}$, and level curves of the objective function.
@@
 5 $1.4833\times 10^{-14}$ $7.5482\times 10^{-16}$
-Example 3.
-Let us consider $F(x)=(x_{1}^{2},x_{2}^{2},x_{1}+x_{2},1),$ the residual function.
-The corresponding least-squares function $\phi(x)$ has a unique global minimizer at $x^{*}=(0,0)^{T}$, which is the only stationary point of $\phi(x)$, $X^{*}=\{x^{*}\}$.
-The rank of
-$J(x)=\left[\begin{array}[]{cc}2x_{1}&0\\ 0&2x_{2}\\ 1&1\\ 0&0\end{array}\right]$
-is 2 for $x\neq x^{*}$ and it is 1 at $x^{*}$.
-Therefore the rank diminishes as we approach $X^{*}$ and (A1) holds everywhere, regardless $L$, except at $x^{*}$.
-Also, since $\|J(x)-J(y)\|\leq 2\|x-y\|$, (A2) is verified.
-The distinguished feature of this example is that the error bound condition (A3) does not hold for every point in a neighborhood of $x^{*}$.
-In fact, by considering the sequence $x_{k}=(1/k,-1/k)^{T}$ it is not hard to show that
-$\lim_{k\to\infty}\frac{\|\nabla\phi(x_{k})\|}{\textup{dist}(x_{k},X^{*})}=\lim_{k\to\infty}\frac{\|\nabla\phi(x_{k})\|}{\|x_{k}\|}=0,$
-and thus (A3) cannot hold.
-Nevertheless, for any $x\in\text{span}\{(1,1)^{T}\}={\cal N}(J(x^{*}))^{\perp}$, from the mean value theorem, we have
-$\|\nabla\phi(x)-\nabla\phi(0)\|=\|\nabla^{2}\phi(\xi)x\|\geq 2\|x\|,$
-where we used the fact that $(1,1)^{T}$ is an eigenvector of $\nabla^{2}\phi(\xi)$ and 2 is a lower bound for the corresponding eigenvalue.
-Therefore, in this example, starting at $x_{0}\in{\cal N}(J(x^{*}))^{\perp}$ and using $L$ such that ${\cal N}(L)={\cal N}(J(x^{*}))^{\perp}$, we force the sequence $\{x_{k}\}$ generated by LMMSS to stay in a subspace where (A3) holds.
-Also, with this choice of $L$, (A1) is satisfied everywhere, and since we have
-$\|J(x)^{T}F(0)\|=0\leq K\|x-0\|^{2},$
-(A4) with $r=1$ is fulfilled and quadratic convergence is achieved, as depicted in Table 3.
-Table 3: Distances of $x_{k}$ to $x^{*}$ produced by LMMSS until the stopping criterion $\|J_{k}^{T}F_{k}\|<10^{-10}$ is achieved for Example 3 with $L=[-1\quad 1]$, for different starting points.
-$k$ $x_{0}^{(1)}=(3,3)^{T}$ $x_{0}^{(2)}=(-2,-2)^{T}$
-1 $2.0097$ $1.2571$
-2 $8.0542\times 10^{-1}$ $3.8494\times 10^{-1}$
-3 $1.5845\times 10^{-1}$ $2.4840\times 10^{-2}$
-4 $1.9403\times 10^{-3}$ $7.6586\times 10^{-6}$
-5 $3.6524\times 10^{-9}$ $2.6331\times 10^{-16}$
-6 $7.4734\times 10^{-17}$
-It is worth mentioning that even though (A3) does not hold for every $x$ in a neighborhood of $x^{*}$ in this example, a weaker Hölderian error bound condition [1] does hold.
-Precisely $\|\nabla\phi(x)\|\geq\|x\|^{3}$ or, in other words, $\nabla\phi$ is Hölder metrically subregular of order 1/3 (see [1, Definition 1]).
+Example 3. (Diminishing rank and (A3)) The residual function
+$F(x)=(\frac{1}{9}\cos x_{1}-x_{2}\sin x_{1},\ \ \frac{1}{9}\sin x_{1}+x_{2}\cos x_{1})^{T}$
+has a set of nonisolated minimizers $X^{*}=\{x\in\mathbb{R}^{2}\mid x_{2}=0\}$, thus $\textup{dist}(x,X^{*})=|x_{2}|$.
+In (behling2019local, , Example 5.3) it was verified that for $x^{*}=(\pi,0)^{T}$, and $\delta=0.1$, (A1) is satisfied, as well as (A2) with $\omega=1$ (because $\|\nabla\phi(x)\|=|x_{2}|$).
+The Jacobian, given by
+$J(x)=\left[\begin{array}[]{cc}-x_{2}\cos x_{1}-\frac{1}{9}\sin x_{1}&\quad-\sin x_{1}\\ \\ -x_{2}\sin x_{1}+\frac{1}{9}\cos x_{1}&\quad\cos x_{1}\end{array}\right]$
+has full rank for $x_{2}\neq 0$ and rank 1 for $x_{2}=0$.
+Thus, the rank of $J(x_{k})$ decreases as $x_{k}$ approaches $X^{*}$ from the outside.
+Moreover, condition (4) is valid in $B(x^{*},\delta)\setminus X^{*}$.
+Unfortunately, (A4) does not hold in this example. But (A3) does, because
+$\|J(x)^{T}F(z)\|$ $=\frac{1}{18}\sqrt{81\sin^{2}(x_{1}-z_{1})+(9x_{2}\cos(x_{1}-z_{1})+\sin(x_{1}-z_{1}))^{2}}$
+$\leq\frac{\sqrt{91}}{81}\|x-z\|,$
+for $x\in B(x^{*},\delta)$ and $z\in X^{*}\cap B(x^{*},\delta)$. Hence, $\sigma=\sqrt{91}/81\approx 0.118$.
+Other constants relevant for the analysis are given approximately by $L_{0}=1.1156$, $L_{4}=0.5612$, $\gamma=0.8949$, $(\sigma_{\min}^{*})^{2}=1.0123$.
+We observe that $\sigma<(\sigma_{\min}^{*})^{2}$, and for $\theta=1.0315$, we ensure $(\sigma_{\min}^{*})^{2}>\theta\sigma$.
+In this case, $\bar{\delta}=0.1=\delta$, and assumptions required by Lemma 6 hold.
+By considering $L=[1\quad 0]$, $\|L\|=1$ and from (42) with $\kappa=1.0001$ we obtain $c_{1}=2.6172$.
+Then, since $\bar{\sigma}=0.1367>\sigma$, we can follow Lemma 14 with $\varepsilon<=0.0016$ and $\eta\in]0.8617,1[$.
+Setting $\eta=0.9$, makes $\varepsilon=0.0016$ in (43) and starting from $x_{0}=(\pi,0.001)^{T}$, and using $\lambda_{k}=\sigma+L_{4}\|x_{k}-\bar{x}_{k}\|$, Theorem 5 applies, and we expect linear convergence, as can be verified in Table 3.
+$k$ $\textup{dist}(x_{k},X^{*})$
+0 0.001
+1 0.00010413
+2 1.0889e-05
+3 1.1392e-06
+4 1.1919e-07
+5 1.247e-08
+6 1.3046e-09
+7 1.365e-10
+8 1.4281e-11
+Table 3: Iterations of LMMSS for Example 3, initial point $x_{0}=(\pi,0.001)^{T}$.
 7 Final remarks
-We have investigated convergence issues of Levenberg-Marquardt method with singular scaling (LMMSS) [5] for nonlinear least-squares problems with non-zero residue.
-Through a couple of modifications in the algorithm proposed in [5, Algorithm 1], mainly, the choice of Levenberg-Marquardt parameter and the line-search, we have established local and global convergence results under mild assumptions.
-It was shown that LMMSS direction is well defined as long as the nullspace of the Jacobian and the nullspace of the singular scaling matrix $L^{T}L$ have trivial intersection (Assumption 1).
-Then, it was demonstrated local superlinear convergence rate under a local error bound condition on the gradient of the least-squares function (Assumption 3) and an assumption on the linearization error of the gradient (Assumption 4) which represents a measure of nonlinearity and residue size.
-These results were established for both the constant rank and diminishing rank cases.
-By using an Armijo line-search and safeguard directions, it was proved that all limit points of the sequence generated by Algorithm 1 are stationary points of the least-squares function.
-We also presented some examples illustrating the assumptions and convergence results.
-In future works we plan to consider Hölder metric subregularity [1] instead of the local error bound used here and address local convergence when a weaker form of Assumption 4 holds, namely when $\|J(x)^{T}F(z)\|={\cal O}(\|x-z\|)$ instead of $\|J(x)^{T}F(z)\|=o(\|x-z\|)$, following the lines of the analysis presented in [2].
-Apart from tackling these questions, we aim to conduct computational experiments to test the algorithm, equipped with suitable stopping criteria, in certain inverse problems to assess its numerical behavior.
+We have investigated the local convergence of Levenberg-Marquardt method with singular scaling (LMMSS) proposed in Boos2024 when applied to nonlinear least-squares problems with nonzero residual.
+Our study reveals that, regardless the Jacobian rank is constant or not in a neighborhood of $x^{*}\in X^{*}$, the convergence rate of $\textup{dist}(x_{k},X^{*})$ to zero depends on a combined measure of nonlinearity and residual size $\|J(x)^{T}F(z)\|$ (see Assumption 3 and 4) and suitable choices of the LM parameter $\lambda_{k}$ (Theorems  25).
+For $\|J(x)^{T}F(z)\|=o(\|x-z\|)$, the local convergence rate is superlinear whereas $\|J(x)^{T}F(z)\|\leq\sigma\|x-z\|$ leads to linear convergence, but only when $\sigma$ is small enough relatively to the error bound constant $\omega$.
+These results are aligned with those of behling2019local but consider the more general case where $L^{T}L$ can be singular, as long as the uniform completeness condition (Definition 1) holds in a suitable neighborhood of $x^{*}\in X^{*}$.
+Furthermore, through a few modifications in the algorithm proposed in (Boos2024, , Algorithm 1), namely, the choice of LM parameter, the full step acceptance criterion and the use of classic LMM direction as a safeguard whenever LMMSS fails to satisfy descent properties; we have established under mild assumptions the global convergence of Algorithm 1: every limit point of the generated sequence is stationary for the nonzero residual nonlinear least squares problem.
+In future works we will look for weaker assumptions to control $\|J(x)^{T}F(z)\|$.
+The condition $\|J(x)^{T}F(z)\|\leq\sigma\|x-z\|$ with small enough $\sigma$ may be too restrictive and hold only in nonzero residual problems with very small residual norm.
+When it does not hold, we cannot even ensure that the “pure” LMMSS converges locally.
+In that case, other methods such as Quasi-Newton or even Newton, when the Hessians of $F_{i}(x)$ are available, might be more appropriate.
 Acknowledgments
-This work was supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360].
-The work of DG was partially supported by CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico), Brazil [grant number 305213/2021-0].
+This work was supported by Brazilian agencies FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360] and CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico) [grant number 305213/2021-0].